### v1

# Disentangled Interpretable Representation for Efficient Long-term Time Series Forecasting

Abstract
Industry 5.0 introduces new challenges for Long-term Time Series Forecasting (LTSF), characterized by high-dimensional, high-resolution data and high-stakes application scenarios. Against this backdrop, developing efficient and interpretable models for LTSF becomes a key challenge. Existing deep learning and linear models often suffer from excessive parameter complexity and lack intuitive interpretability. To address these issues, we propose DiPE-Linear, a Disentangled interpretable Parameter-Efficient Linear network. DiPE-Linear incorporates three temporal components: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM). These components alternate between learning in the frequency and time domains to achieve disentangled interpretability. The decomposed model structure reduces parameter complexity from quadratic in fully connected networks (FCs) to linear and computational complexity from quadratic to log-linear. Additionally, a Low-Rank Weight Sharing policy enhances the model’s ability to handle multivariate series. Despite operating within a subspace of FCs with limited expressive capacity, DiPE-Linear demonstrates comparable or superior performance to both FCs and nonlinear models across multiple open-source and real-world LTSF datasets, validating the effectiveness of its sophisticatedly designed structure. The combination of efficiency, accuracy, and interpretability makes DiPE-Linear a strong candidate for advancing LTSF in both research and real-world applications. The source code is available at https://github.com/wintertee/DiPE-Linear.
Index Terms:
Time series analysis, Time series forecasting, Efficient neural network, Interpretable artificial intelligenceI Introduction
Long-term Time Series Forecasting (LTSF) is of significant importance in numerous domains, particularly in industrial applications such as cloud computing [1, 2], energy optimization [3, 4], and manufacturing [5, 6]. Recent advancements in deep learning, including Recurrent Neural Networks (RNNs) [7], Convolutional Neural Networks (CNNs) [8, 9, 10, 11], Transformers [12, 13, 14, 15, 16, 17, 18, 19, 20], Multilayer Perceptrons (MLPs)[21, 22, 23, 24, 25, 26] and Fully Connected linear models (FCs) [27, 28, 29] have demonstrated promising performance for LTSF tasks. With the advent of Industry 5.0 and the era of big data, time-series data processing faces three primary challenges: high dimensionality, high resolution, and regulatory industrial scenarios. As regulatory emphasis on carbon footprint and trustworthiness increases[30], developing an efficient and interpretable time-series forecasting method has emerged as a critical challenge.
Long-term forecasting for high-resolution time-series data remains challenging. Metrics like CPU load and network traffic in cloud environments are sampled at minute intervals, yet predictions must span days to support proactive resource allocation [31, 32, 20]. Similarly, advanced sensors in industrial settings generate vast minute-level data, requiring accurate long-term forecasting for equipment monitoring and operational optimization [33]. Existing methods face key limitations: RNNs struggle with parallelization [7], CNNs (e.g., TimesNet [11]) impose high computational costs, and Transformers, despite linear complexity improvements [16, 17], often overfit on long input sequences [27]. Fully connected models offer simplicity and competitive performance but suffer from quadratic complexity and parameter redundancy (Figs. 1(a), 1(b) and 1(c)). This raises the pivotal question: How can a model achieve parameter efficiency to effectively handle the challenges of long-term forecasting for high-resolution data?
In high-stakes industrial scenarios, interpretability are indispensable requirements for achieving trustworthy AI [34, 35]. In this context, domain experts must understand the model to trust its predictions. While leading deep learning models offer impressive precision, their black-box nature significantly undermines user trust. Fully connected linear models, by contrast, provide better interpretability by design. However, it remains challenging to fully and directly understand the features encoded by these models through their two-dimensional weight matrices, whether represented in the time domain (Figs. 1(a) and 1(b)) or frequency domain (Fig. 1(c)). This introduces the second key question: How can a model ensure interpretability, enabling domain experts to understand and trust its predictions in high-stakes industrial applications?
To address these challenges, we revisit the performance of FCs in LTSF, seeking inductive biases to reduce complexity and enhance interpretability. Figs. 1(a) and 1(b) reveals that highly correlated weights along the diagonal suggest parameter redundancy in the time domain, while the high noise in off-diagonal regions in the frequency domain as shown in Fig. 1(c) indicates the independence of frequency components. This motivates our frequency independence hypothesis. Furthermore, time series often exhibit varying contributions of input time points to future predictions; for example, the most recent data typically plays a greater role in predicting future values. Thus, modeling the temporal significance of input sequences is crucial. Similarly, different frequency components contribute unequally[36], with significant frequencies often possessing higher signal-to-noise ratios and deserving more model attention. Finally, in multivariate time series, variable dependencies must be captured in an explainable manner, as variables may either share patterns or act independently.
Based on the aforementioned motivations, we propose a novel Disentangled interpretable and Parameter-Efficient Linear model (DiPE-Linear) for long-term time series forecasting. As a deep linear model, it is structured into three interpretable linear sub-modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM). The SFA module extracts and enhances prominent frequencies in the frequency domain, and the STA module selects input time points in the time domain that are strongly correlated with the predicted sequence. Meanwhile, the IFM module assumes independence across different frequencies and directly maps the input series to the output series at each frequency component in frequency domain. By leveraging disentangled representations in both the time and frequency domains, the model achieves reduced complexity with linear parameter growth and log-linear computational scaling. Additionally, as shown in Fig. 1(d), these representations enhance domain experts’ trust by providing clear insights into the learned features and the model’s reasoning process.
These sub-modules employ a low-rank weight-sharing policy to efficiently model inter-variate relationships, reducing model complexity while maintaining strong performance. We further introduce a novel loss function, SFALoss, which is defined as a sum of time-domain loss and frequency-domain loss weighted by frequency-domain attention in the SFA module, guiding the model to prioritize significant frequencies with less noise.
Additionally, by avoiding non-linearity within these modules, our model remains a linear model while achieving comparable or SOTA performance in LSTF with significantly fewer parameters, as shown in Fig. 2. This lightweight design enhances model’s interpretability and establishes a more reliable and efficient prediction pipeline, offering valuable insights for future research.
In summary, our contributions are multifold:
We propose DiPE-Linear, a deep linear model that achieves efficiency, significantly reducing parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear.
Disentangled representation of model components in both time and frequency domain offers superior interpretability, thereby enhancing its trustworthiness for industrial applications.
Extensive experiments demonstrate that our method outperforms existing approaches, offering superior performance, efficiency, and interpretability.
II Related Work
II-A Efficient Architectures for Long-Term Forecasting
Long-term time series forecasting (LTSF) requires models capable of efficiently handling long sequences. Early Transformer-based models [12, 13, 14, 15, 16, 17] progressively reduced computational complexity from to . However, due to the inherently complex structure of Transformers, these models still contain a large number of parameters and are highly prone to overfitting when dealing with excessively long input sequences in LTSF tasks[27].
DLinear[27] was the first to propose using a simple single-layer fully connected network to replace the complex Transformer structure. Although the fully connected layer has a computational complexity of , its simple structure results in significantly lower complexity compared to other Transformer-based models. Nevertheless, DLinear still suffers from notable parameter redundancy. FITS[29] addressed this by employing low-pass filters to discard high-frequency information in the frequency domain, drastically reducing the number of parameters in the fully connected network to the scale of 10k. SparseTSF[37], on the other hand, introduced Cross-Period Sparse Forecasting, which achieves a parameter count on the scale of 1k through downsampling and cross-period parameter sharing. However, both FITS and SparseTSF sacrifice accuracy to achieve parameter reduction[38, 37]. Further exploration into addressing parameter redundancy in fully connected models remains an open research question.
II-B Exploiting Frequency-Domain Characteristics
Recent advancements in time series forecasting have increasingly leveraged frequency-domain knowledge to improve performance. One prominent approach incorporates frequency-domain information as auxiliary features to enhance the modeling of long-term dependencies. For example, FEDformer[17] introduces a DFT-based frequency-enhanced attention mechanism, where attention weights are computed from the spectra of queries and keys, with the weighted sum calculated directly in the frequency domain. Similarly, FiLM[39] applies Fourier analysis to retain historical information while filtering out noise. Alternatively, some methods operate entirely in the frequency domain. FreTS[40] employs a frequency-domain MLP to capture both channel-wise and temporal dependencies, while FITS[29] adopts a fully connected network specifically designed for spectral representations.
Another promising direction incorporates frequency-domain loss functions into time-domain training objectives to capture spectral patterns more effectively. For example, FTMixer [26] and FreDF [41] add frequency-domain loss components on top of time-domain losses, enhancing the model’s ability to align with the ground truth in the frequency domain.
II-C Channel-Independence for Multivariate Time Series
DLinear[27] was the first to introduce the Channel-Independent strategy for multivariate time series forecasting (the terms ”channel” and ”variate” are used interchangeably throughout this paper for simplicity). Although this approach ignores causal relationships between variables, it achieves significant performance improvements. PatchTST[19] further validated this idea in Transformer-based models, demonstrating that independently processing variables while sharing the same model weights across variables can be highly effective.
However, RLinear[28] shows that for variables with distinct periodic characteristics, using independent weights for each variable can further enhance performance. Subsequent research has focused on finding an optimal balance between weight sharing and weight independence. For instance, MLinear[42] dynamically adjusts between Weight Sharing and Weight Independent strategies based on the temporal semantics of different sequences, while CSC[43] reassigns weight-variable mappings during the validation phase to achieve variable clustering. Nevertheless, these methods come with limitations. MLinear requires calculating both Weight Sharing and Weight Independent outputs before combining them, which increases computational overhead. CSC, on the other hand, relies on a strong prior assumption that variables clustered together should share the same weights, which may not hold universally.
III Method
DiPE-Linear is constructed from several essential components. We begin by outlining the preliminaries related to TSF and the Fourier Transform, providing the foundation for our approach. Next, we present the architecture of DiPE-Linear, detailing its three core modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM), under assumption that the input series is univariate. Then, we describe how the Low-rank Weight Sharing method generalizes those modules to multivariate time series. Lastly, we introduce SFALoss, a novel loss function that combines weighted losses in both the frequency and time domains, guiding the model to prioritize informative patterns while enhancing predictive performance.
III-A Preliminaries
III-A1 Problem Definition
The multivariate TSF problem can be formally defined as follows: Given a historical time series , where is the number of variates and is the look-back length of input time series, the objective is to predict the future time series , where is the forecasting horizon.
III-A2 Discrete Fourier Transform
Time series data can be viewed as discrete samples of a continuous signal over time. In many signal processing and time-series forecasting applications, the Discrete Fourier Transform (DFT) is a commonly used method to analyze the frequency components of such discrete signals. For a discrete-time sequence of length , the DFT transforms the time-domain signal into its frequency-domain representation. This transformation is particularly useful for identifying underlying periodic patterns or extracting frequency components that contribute to the signal.
The DFT of a time series is defined as:
Once the signal is represented in the frequency domain, various operations, such as filtering or modifying specific frequency components, can be applied. The inverse Discrete Fourier Transform (iDFT) allows for transforming the modified frequency-domain signal back to the time domain. The iDFT is expressed as:
This transformation provides the flexibility to manipulate signals in the frequency domain and revert them back to the time domain for further analysis or application. The ability to move between domains seamlessly is essential for efficient signal processing.
III-A3 Duality Between Time and Frequency Domains
The DFT exhibits a fundamental property known as duality between the time and frequency domains. This duality manifests as convolution in the time domain being equivalent to multiplication in the frequency domain, and vice versa. For two time-domain signals, and , their convolution:
has a frequency-domain equivalent:
Pointwise multiplication in the time domain, on the other hand, corresponds to convolution in the frequency domain. This relationship simplifies many operations, particularly in tasks where convolution is computationally intensive in the time domain.
Thus, the DFT provides a powerful tool for analyzing time series in the frequency domain, leveraging the duality between time and frequency domains to simplify complex operations such as convolution. The ability to transition between domains efficiently enables advanced signal processing techniques for both single-channel and multichannel data.
III-B Static Frequential Attention
Extracting frequencies is a common practice, as previous FITS [29] employed a low-pass filter to eliminate high-frequency signals, while FedFormer [17] randomly selects a subset of frequencies as key features. However, such strong prior assumptions may lead the model to discard useful information. Therefore, we aim for the model to learn which frequency components to focus on.
Specifically, to ensure the model’s capability to select and amplify significant frequencies, we introduce SFA module as a frequency-domain filter. Additionally, to preserve the temporal structure of the series and avoid interference with subsequent temporal feature extraction, we constrained this filter to be a zero-phase filter. The SFA operates by first applying real Fast Fourier Transform (rFFT) to transform the time-domain data into the frequency domain. In this domain, a learned filter is applied via element-wise multiplication, selectively amplifying or suppressing specific frequency components. As a zero-phase filter, it exclusively modifies the amplitude of the signal while preserving the original phase information. Finally, inverse rFFT (irFFT) is used to reverse the filtered data back to the time domain. Specifically, the enhanced signal is given by
where represents the input univariate series, denotes the rFFT, is the irFFT, is the learnable static frequential attention map, and denotes element-wise multiplication.
III-C Static Temporal Attention
Similarly, to enable the model to have the capability of capturing important input time points in the temporal domain, we introduce the STA as a second component of our model. This module enables the input time series to undergo element-wise multiplication with a learned temporal attention map, effectively assigning appropriate importance to relevant historical time points and thereby enhancing the model’s ability to capture temporal dependencies. The processed signal is thus given by
where is the learnable time-domain attention map.
III-D Independent Frequential Mapping
While both the SFA and STA modules function as part of the input series pre-processing, the IFM module directly maps the input historical series to the forecasted output. To process the time series as a whole, leveraging the spectrum is a natural approach, as it encapsulates all the information about the entire sequence and enables better handling of periodic components. Properly and efficiently processing the spectrum remains a key challenge. Fig. 1(c) illustrates that the two-dimensional matrix used by FITS [29] to handle the spectrum is sparse, with only the diagonal weights being significant. This suggests that, for a time series, there is relatively strong independence between different frequencies, allowing each frequency to be handled separately.
We thus treat each frequency independently for two key reasons: (1) Apart from harmonic frequencies, there is generally little correlation observed between different frequencies in time series data.[41] (2) Although harmonic frequencies tend to exhibit strong correlations, we avoid explicitly modeling these inter-frequency dependencies to mitigate potential multicollinearity in linear regression.
Therefore, for each frequency component, we apply a complex-valued multiply-accumulate operation independently. To transform the input space to the output space , we apply zero-padding to the input series, extending its length to before performing the rFFT. At the end, the last values are extracted and used as the predicted output. The forecasted output is thus defined as:
where is the complex-valued weight and is the complex-valued bias.
This process can be equivalently described as a 1D convolution in time domain, as explained in equation 3 and 4:
where is the convolution kernel in time domain. The extremely large convolution kernel distinguishes it from traditional CNNs that use much smaller kernels. This larger kernel allows each output to capture a global receptive field, enabling the model to incorporate information from the entire input sequence.
III-E Low-rank Weight Sharing
Weight sharing is a common practice in time series forecasting models [27, 29, 19], based on the assumption that different channels of a time series exhibit similar patterns. However, this assumption often does not hold. For example, in Weather datasets, variables such as temperature, pressure, and rainfall may follow distinct patterns. In such cases, training weights independently for each channel can improve performance [28], but this approach significantly increases the number of parameters by a factor of , making the model complexity significantly higher.
To address this, we propose a more efficient alternative. By clustering variables based on their pattern similarities, we can reduce the number of weight sets, balancing the trade-off between model efficiency and performance. This clustering approach not only enhances computational efficiency by reducing the number of learned parameters but also improves accuracy, as the shared weights capture more robust patterns. Specifically, well-structured clustering ensures that the learned representations generalize better across similar variables. Inspired by mixture-of-experts[44] and dynamic convolution[45], we introduce a novel Low-rank Weight Sharing architecture, as shown in Fig. 4, leveraging these clustered patterns to optimize both parameter efficiency and predictive performance.
Specifically, for a multivariate time series with variables, we introduce a hyperparameter , where denotes the number of independent weight sets to be learned, with . For each component, distinct sets of weights are learned, alongside a static routing matrix. The low-rank routing matrix is designed to linearly combine the weights and assign each variable to an appropriate weight set. We regularize the static routing matrix using the Softmax function, where is the Softmax temperature:
We denote as the set of mappings representing any given module, corresponding to the independent weight sets learned. For variate , the weight actually applied to channel is:
III-F SFALoss
Building upon the insights of FreDF [41], we introduce SFALoss (denoted as ), which combines Weighted Mean Absolute Error (WMAE) in the frequency domain () and MSE in the time domain (). For the frequency domain component (), our SFA module selectively subtracts key frequencies while attenuating noisy frequencies. To refine this process, we modulate the frequency-domain loss through an element-wise multiplication with the SFA weighting factor, , effectively filtering out loss contributions from less relevant frequencies. The WMAE loss in the frequency domain is formulated as follows:
where and are respectively ground truth and predicted future time series in frequency-domain. While computing the loss, is detached from the computational graph, ensuring that no gradients are backpropagated through it.
We also preserve the MSE loss in time domain :
Finally, the overall SFAloss function is defined as follow:
where is hyperparameter that balances the contribution of the two loss terms, allowing us to achieve a trade-off between them.
IV Experiment
IV-A Experimental Setup
IV-A1 Dataset
To comprehensively evaluate the performance and generalization capability of our model, we selected a diverse set of eight datasets, including six for LTSF and two for short-term time series forecasting (STSF). These datasets come from various domains, with different sizes and sampling frequencies, offering a robust benchmark for assessing the versatility and scalability of our approach across a wide range of forecasting tasks. Details are shown in Table I.
For the LTSF task, we utilized six public datasets:
The ETT111https://github.com/zhouhaoyi/ETDataset dataset, including ETTh1, ETTh2, ETTm1 and ETTm2, comprises load and oil temperature data from two power transformers at different stations.
The Electricity222https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 dataset includes electricity consumption data, measured in kilowatt-hours (kWh), from 321 clients, spanning from 2012 to 2014.
The Weather333https://www.bgc-jena.mpg.de/wetter/ dataset consists of 21 meteorological indicators, sampled every 10 minutes, covering the entire year of 2020.
Moreover, to validate the effectiveness of the model in real-world production environments, we selected two workload dataset collected from ByteDance’s cloud servers, serveing as private LTSF datasets:
The FaaS includes 15-day query-per-second (QPS) data from the public cloud platform.
The IaaS includes 45-day CPU usage data from the private cloud platform.
To further demonstrate the generalizability of our model, we extended our evaluation to two short-term forecasting tasks using the following datasets:
The Illness444https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html dataset records weekly influenza-like illness cases as reported by the Centers for Disease Control and Prevention (CDC) in the United States, from 2002 to 2021.
The M5555https://www.kaggle.com/competitions/m5-forecasting-accuracy/ dataset contains hierarchical sales data from Walmart, covering stores in three U.S. states.
| Category | Dataset | Domain | Variates | Samples | Interval |
| LTSF | ETTh1 | Industry | 7 | 17420 | 1 hour |
| ETTh2 | Industry | 7 | 17420 | 1 hour | |
| ETTm1 | Industry | 7 | 69680 | 15 min | |
| ETTm2 | Industry | 7 | 69680 | 15 min | |
| Electricity | Energy | 321 | 26304 | 1 hour | |
| Weather | Environment | 21 | 52696 | 1 hour | |
| FaaS | Cloud Computing | 1014 | 23041 | 1 min | |
| IaaS | Cloud Computing | 168 | 69045 | 1 min | |
| STSF | Illness | Medical | 7 | 966 | 1 week |
| M5 | Market | 30 | 1941 | 1 day |
| DiPE-Linear | FITS | RLinear | DLinear | Nlinear | |||||||
| (Ours) | (ICLR-24) | (arXiv-23) | (AAAI-23) | (AAAI-23) | |||||||
| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| ETTh1 | 96 | 0.3690.000 | 0.3930.000 | 0.3800.000 | 0.4030.000 | 0.3760.004 | 0.4010.004 | 0.3880.010 | 0.4140.012 | 0.3840.004 | 0.4050.003 |
| 192 | 0.4070.000 | 0.4150.000 | 0.4170.001 | 0.4250.001 | 0.4160.002 | 0.4260.002 | 0.4250.005 | 0.4360.005 | 0.4150.003 | 0.4240.002 | |
| 336 | 0.4240.000 | 0.4270.000 | 0.4360.000 | 0.4400.000 | 0.4440.002 | 0.4440.002 | 0.4690.010 | 0.4690.009 | 0.4450.002 | 0.4430.002 | |
| 720 | 0.4090.000 | 0.4390.000 | 0.4320.000 | 0.4560.000 | 0.4740.002 | 0.4810.002 | 0.5300.021 | 0.5330.015 | 0.4390.002 | 0.4700.000 | |
| ETTh2 | 96 | 0.2750.001 | 0.3360.001 | 0.2720.000 | 0.3360.000 | 0.2700.001 | 0.3350.001 | 0.2800.005 | 0.3450.004 | 0.2760.002 | 0.3380.001 |
| 192 | 0.3250.001 | 0.3720.000 | 0.3310.000 | 0.3740.000 | 0.3350.004 | 0.3800.002 | 0.3580.014 | 0.3990.011 | 0.3450.004 | 0.3820.002 | |
| 336 | 0.3500.002 | 0.3930.001 | 0.3540.000 | 0.3950.000 | 0.3660.005 | 0.4090.003 | 0.4390.017 | 0.4570.009 | 0.3750.009 | 0.4110.004 | |
| 720 | 0.3750.002 | 0.4150.001 | 0.3780.000 | 0.4230.000 | 0.4140.003 | 0.4470.001 | 0.6570.062 | 0.5730.026 | 0.4080.008 | 0.4460.003 | |
| ETTm1 | 96 | 0.3090.000 | 0.3500.000 | 0.3090.000 | 0.3520.001 | 0.3090.001 | 0.3520.001 | 0.3120.005 | 0.3580.006 | 0.3180.007 | 0.3570.005 |
| 192 | 0.3390.000 | 0.3690.000 | 0.3390.001 | 0.3690.001 | 0.3410.004 | 0.3700.003 | 0.3500.006 | 0.3860.008 | 0.3500.007 | 0.3770.005 | |
| 336 | 0.3670.000 | 0.3860.000 | 0.3680.001 | 0.3860.001 | 0.3690.003 | 0.3870.003 | 0.3790.005 | 0.4030.006 | 0.3780.005 | 0.3930.004 | |
| 720 | 0.4160.000 | 0.4130.000 | 0.4160.000 | 0.4130.000 | 0.4150.001 | 0.4110.001 | 0.4390.011 | 0.4430.012 | 0.4200.003 | 0.4150.002 | |
| ETTm2 | 96 | 0.1620.000 | 0.2520.000 | 0.1620.000 | 0.2530.000 | 0.1620.000 | 0.2510.000 | 0.1630.001 | 0.2550.001 | 0.1620.000 | 0.2520.001 |
| 192 | 0.2160.000 | 0.2890.000 | 0.2160.000 | 0.2910.000 | 0.2160.000 | 0.2900.000 | 0.2190.002 | 0.2970.001 | 0.2160.000 | 0.2910.000 | |
| 336 | 0.2680.000 | 0.3240.000 | 0.2680.000 | 0.3260.000 | 0.2680.001 | 0.3260.000 | 0.2720.003 | 0.3340.003 | 0.2680.000 | 0.3260.000 | |
| 720 | 0.3530.000 | 0.3790.000 | 0.3490.000 | 0.3780.000 | 0.3540.001 | 0.3840.000 | 0.3670.004 | 0.3980.003 | 0.3490.000 | 0.3790.000 | |
| Electricity | 96 | 0.1320.000 | 0.2280.001 | 0.1340.000 | 0.2310.000 | 0.1350.000 | 0.2320.000 | 0.1330.000 | 0.2290.000 | 0.1330.000 | 0.2280.000 |
| 192 | 0.1480.000 | 0.2450.000 | 0.1490.000 | 0.2440.000 | 0.1500.000 | 0.2460.000 | 0.1480.000 | 0.2430.000 | 0.1480.000 | 0.2420.000 | |
| 336 | 0.1620.000 | 0.2610.000 | 0.1650.000 | 0.2600.000 | 0.1660.000 | 0.2620.000 | 0.1620.000 | 0.2610.000 | 0.1640.000 | 0.2580.000 | |
| 720 | 0.1980.002 | 0.2960.001 | 0.2040.000 | 0.2930.000 | 0.2060.000 | 0.2940.000 | 0.1960.000 | 0.2930.001 | 0.2030.000 | 0.2910.000 | |
| Weather | 96 | 0.1420.001 | 0.2010.001 | 0.1420.000 | 0.1920.000 | 0.1430.000 | 0.1940.000 | 0.1420.000 | 0.2020.000 | 0.1420.001 | 0.1920.001 |
| 192 | 0.1870.003 | 0.2530.002 | 0.1850.000 | 0.2340.000 | 0.1860.000 | 0.2350.000 | 0.1840.001 | 0.2490.002 | 0.1830.000 | 0.2340.000 | |
| 336 | 0.2340.001 | 0.2930.002 | 0.2350.000 | 0.2760.000 | 0.2360.000 | 0.2750.000 | 0.2360.001 | 0.2930.001 | 0.2340.001 | 0.2760.000 | |
| 720 | 0.3060.004 | 0.3480.003 | 0.3070.000 | 0.3280.000 | 0.3080.000 | 0.3270.000 | 0.3060.002 | 0.3490.002 | 0.3080.000 | 0.3300.000 | |
| FaaS | 96 | 0.2800.000 | 0.2510.000 | 0.3090.000 | 0.2660.000 | 0.3060.000 | 0.2640.000 | 0.3030.000 | 0.2640.000 | 0.3050.000 | 0.2660.000 |
| 192 | 0.3140.000 | 0.2790.000 | 0.3420.000 | 0.2930.000 | 0.3440.000 | 0.2920.000 | 0.3390.000 | 0.2940.000 | 0.3400.000 | 0.2920.000 | |
| 336 | 0.3510.000 | 0.3090.000 | 0.3680.000 | 0.3160.000 | 0.3750.000 | 0.3190.000 | 0.3670.000 | 0.3230.000 | 0.3640.000 | 0.3140.000 | |
| 720 | 0.3790.000 | 0.3320.000 | 0.4150.000 | 0.3480.000 | 0.4360.000 | 0.3500.000 | 0.4170.000 | 0.3640.001 | 0.4100.000 | 0.3440.000 | |
| IaaS | 96 | 0.7890.000 | 0.6670.000 | 0.7990.000 | 0.7020.000 | 0.7970.000 | 0.6880.000 | 0.7950.000 | 0.6990.000 | 0.7990.000 | 0.6820.000 |
| 192 | 0.8170.000 | 0.6980.000 | 0.8390.000 | 0.7230.000 | 0.8310.000 | 0.7000.000 | 0.8230.000 | 0.7170.000 | 0.8200.000 | 0.7020.000 | |
| 336 | 1.1800.015 | 0.8390.010 | 0.9980.007 | 0.8550.005 | 1.2550.012 | 0.8220.007 | 0.8420.000 | 0.7310.000 | 0.8980.000 | 0.7710.000 | |
| 720 | 0.8690.000 | 0.7360.000 | 0.8990.000 | 0.8390.000 | 0.9380.000 | 0.8280.000 | 0.8810.000 | 0.7410.000 | 0.9240.000 | 0.7850.000 | |
| Illness | 24 | 2.2600.000 | 0.9510.000 | 2.2890.000 | 0.9750.000 | 2.2580.000 | 0.9300.000 | 2.4670.000 | 1.0790.000 | 2.2880.000 | 0.9620.000 |
| 36 | 2.1620.000 | 0.9400.000 | 2.2920.000 | 0.9860.000 | 2.2130.000 | 0.9430.000 | 2.5100.000 | 1.0890.000 | 2.2210.000 | 0.9640.000 | |
| 48 | 2.1840.000 | 0.9560.000 | 2.2660.000 | 0.9920.000 | 2.3090.000 | 0.9790.000 | 2.5100.000 | 1.0920.000 | 2.1770.000 | 0.9700.000 | |
| 60 | 2.0530.000 | 0.9570.000 | 2.2210.000 | 0.9960.000 | 2.1830.000 | 0.9520.000 | 2.5990.000 | 1.1210.000 | 2.2160.000 | 0.9850.000 | |
| M5 | 24 | 0.4900.000 | 0.4970.000 | 0.4990.000 | 0.5020.000 | 0.5000.000 | 0.5030.000 | 0.5030.000 | 0.5050.000 | 0.4980.000 | 0.5020.000 |
| 36 | 0.5150.000 | 0.5110.000 | 0.5230.000 | 0.5290.000 | 0.5240.000 | 0.5170.000 | 0.5330.000 | 0.5210.000 | 0.5220.000 | 0.5150.000 | |
| 48 | 0.5420.000 | 0.5250.000 | 0.5500.000 | 0.5300.000 | 0.5520.000 | 0.5310.000 | 0.5630.000 | 0.5370.000 | 0.5490.000 | 0.5290.000 | |
| 60 | 0.5650.000 | 0.5380.000 | 0.5720.000 | 0.5420.000 | 0.5750.000 | 0.5430.000 | 0.5810.000 | 0.5470.000 | 0.5700.000 | 0.5410.000 | |
| Best count | 32 | 26 | 7 | 5 | 7 | 7 | 6 | 1 | 9 | 6 | |
| DiPE-Linear | iTransformer | PatchTST† | TimesNet† | Crossformer† | MICN† | FEDformer† | |||||||||
| (Ours) | (ICLR-24) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICML-22) | |||||||||
| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
| ETTh1 | 96 | 0.369 | 0.393 | 0.386 | 0.405 | 0.370 | 0.399 | 0.384 | 0.402 | 0.386 | 0.429 | 0.396 | 0.427 | 0.376 | 0.415 |
| 192 | 0.407 | 0.415 | 0.424 | 0.440 | 0.413 | 0.421 | 0.557 | 0.436 | 0.419 | 0.444 | 0.430 | 0.453 | 0.423 | 0.446 | |
| 336 | 0.424 | 0.427 | 0.449 | 0.460 | 0.422 | 0.436 | 0.491 | 0.469 | 0.440 | 0.461 | 0.433 | 0.458 | 0.444 | 0.462 | |
| 720 | 0.409 | 0.439 | 0.495 | 0.487 | 0.447 | 0.466 | 0.521 | 0.500 | 0.519 | 0.524 | 0.474 | 0.508 | 0.469 | 0.492 | |
| ETTh2 | 96 | 0.275 | 0.336 | 0.297 | 0.348 | 0.274 | 0.336 | 0.340 | 0.374 | 0.276 | 0.338 | 0.289 | 0.357 | 0.332 | 0.374 |
| 192 | 0.325 | 0.372 | 0.371 | 0.403 | 0.339 | 0.379 | 0.402 | 0.414 | 0.345 | 0.382 | 0.409 | 0.438 | 0.407 | 0.446 | |
| 336 | 0.350 | 0.393 | 0.404 | 0.428 | 0.329 | 0.380 | 0.452 | 0.452 | 0.375 | 0.411 | 0.417 | 0.452 | 0.400 | 0.447 | |
| 720 | 0.375 | 0.415 | 0.424 | 0.444 | 0.379 | 0.422 | 0.462 | 0.468 | 0.408 | 0.446 | 0.426 | 0.473 | 0.412 | 0.469 | |
| ETTm1 | 96 | 0.309 | 0.350 | 0.300 | 0.353 | 0.290 | 0.342 | 0.338 | 0.375 | 0.316 | 0.373 | 0.314 | 0.360 | 0.326 | 0.390 |
| 192 | 0.339 | 0.369 | 0.345 | 0.382 | 0.332 | 0.369 | 0.371 | 0.387 | 0.377 | 0.411 | 0.359 | 0.387 | 0.365 | 0.415 | |
| 336 | 0.367 | 0.386 | 0.374 | 0.398 | 0.369 | 0.392 | 0.410 | 0.411 | 0.431 | 0.442 | 0.398 | 0.413 | 0.392 | 0.425 | |
| 720 | 0.416 | 0.413 | 0.429 | 0.430 | 0.416 | 0.420 | 0.478 | 0.450 | 0.600 | 0.547 | 0.459 | 0.464 | 0.446 | 0.458 | |
| ETTm2 | 96 | 0.162 | 0.252 | 0.175 | 0.266 | 0.165 | 0.255 | 0.187 | 0.267 | 0.421 | 0.461 | 0.178 | 0.273 | 0.180 | 0.271 |
| 192 | 0.216 | 0.289 | 0.242 | 0.312 | 0.220 | 0.292 | 0.249 | 0.309 | 0.503 | 0.519 | 0.245 | 0.316 | 0.252 | 0.318 | |
| 336 | 0.268 | 0.324 | 0.282 | 0.340 | 0.274 | 0.329 | 0.321 | 0.351 | 0.611 | 0.580 | 0.295 | 0.350 | 0.324 | 0.364 | |
| 720 | 0.353 | 0.379 | 0.378 | 0.398 | 0.362 | 0.385 | 0.497 | 0.403 | 0.996 | 0.750 | 0.389 | 0.406 | 0.410 | 0.420 | |
| ECL | 96 | 0.132 | 0.228 | 0.132 | 0.229 | 0.129 | 0.222 | 0.168 | 0.272 | 0.187 | 0.283 | 0.159 | 0.267 | 0.186 | 0.302 |
| 192 | 0.148 | 0.245 | 0.151 | 0.246 | 0.147 | 0.240 | 0.184 | 0.289 | 0.258 | 0.330 | 0.168 | 0.279 | 0.197 | 0.311 | |
| 336 | 0.162 | 0.261 | 0.167 | 0.264 | 0.163 | 0.259 | 0.198 | 0.300 | 0.323 | 0.369 | 0.196 | 0.308 | 0.213 | 0.328 | |
| 720 | 0.198 | 0.296 | 0.194 | 0.286 | 0.197 | 0.290 | 0.220 | 0.320 | 0.404 | 0.423 | 0.203 | 0.312 | 0.233 | 0.344 | |
| Weather | 96 | 0.142 | 0.201 | 0.159 | 0.208 | 0.149 | 0.198 | 0.172 | 0.220 | 0.153 | 0.217 | 0.161 | 0.226 | 0.238 | 0.314 |
| 192 | 0.187 | 0.253 | 0.200 | 0.248 | 0.194 | 0.241 | 0.219 | 0.261 | 0.197 | 0.269 | 0.220 | 0.283 | 0.275 | 0.329 | |
| 336 | 0.234 | 0.293 | 0.253 | 0.289 | 0.245 | 0.282 | 0.280 | 0.306 | 0.252 | 0.311 | 0.275 | 0.328 | 0.339 | 0.377 | |
| 720 | 0.306 | 0.348 | 0.321 | 0.338 | 0.314 | 0.334 | 0.365 | 0.359 | 0.318 | 0.363 | 0.311 | 0.356 | 0.389 | 0.409 | |
| Best count | 16 | 14 | 1 | 1 | 8 | 11 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
IV-A2 Implementation details
All experiments are implemented in PyTorch [46] on an NVIDIA 4090D GPU. The models are trained using the Adam optimizer [47] with a learning rate of for epochs. Batch sizes are set to 64 and 8 for LTSF and STSF datasets, respectively. For Electricity and Weather datasets, , and otherwise. Input sequence lengths are , , and for public LTSF, private LTSF, and STSF datasets, respectively. Following [45], we apply linear annealing for the softmax temperature in Low-rank Weight Sharing to ensure balanced learning during early epochs.
IV-A3 Baselines
In this work, we focus on evaluating and comparing linear models, benchmarking our approach against leading methods such as FITS [29], RLinear [28], DLinear, and NLinear [27]. To further demonstrate the effectiveness of our model, we also compare it with nonlinear deep learning models, including iTransformer [20], PatchTST [48], TimesNet [11], Crossformer [18], MICN [10], and FEDformer [17], showing that our model achieves state-of-the-art or comparable performance across both linear and nonlinear baselines. Additionally, SparseTSF [37] and FITS [29] are selected as parameter-efficient baselines to highlight the efficiency of our model design.
IV-A4 Evaluation Metrics
We follow the established practices in prior works to evaluate forecasting performance using two primary metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE).
IV-B Main Results
IV-B1 Versus FCs
We first compared our model DiPE-Linear with other leading FCs. To ensure a fair evaluation, all models were trained using the same trainer settings. Throughout the training process, we carefully monitored each model to ensure convergence under these consistent conditions. All models use the optimal model hyperparameters provided in the authors’ original implementations. Table II presents the evaluation results for the linear models on the LTSF and STSF datasets.
Across all eight datasets, DiPE-Linear consistently delivers comparable or even superior performance. On datasets with smaller sample sizes (ETTh1, ETTh2, Faas, Iaas, Illness, and M5), our model significantly outperforms its counterparts, demonstrating its ability to effectively capture temporal patterns with limited training data while avoiding overfitting. On larger datasets (ETTm1, ETTm2, Electricity, and Weather), DiPE-Linear achieves results comparable to other linear models.
One notable advantage of DiPE-Linear is its substantially reduced parameter count compared to FCs. By operating within a low-dimensional manifold in the FCs’ model space, DiPE-Linear achieves performance comparable to or better than FCs, suggesting that our carefully designed model space contains the optimal solution to TSF tasks. This highlights the robustness and effectiveness of our architecture in representing the essential solution space of FCs with minimal complexity.
IV-B2 Versus nonlinear models
Next, we compare DiPE-Linear with several nonlinear models, including Transformer-based models amd CNN-based models. Table III presents the evaluation results for these nonlinear models on public LTSF datasets. For private LTSF datasets, excessively long input lengths result in high memory usage for nonlinear models, leading to unacceptably high deployment costs in production environments. For STSF datasets, the limited amount of training data makes the models highly prone to overfitting, making training challenging. Therefore, we do not report model performance on these two types of datasets. To ensure a fair comparison, all nonlinear models were trained with input lengths , and the best results were selected for evaluation.
Our model demonstrates a slight performance advantage over PatchTST, while requiring only times the parameter count. Furthermore, our model significantly outperforms other nonlinear models, highlighting its competitive advantages in both efficiency and performance.
IV-B3 Parameter complexity and efficiency
Recent studies have focused on parameter-efficient models, and we selected SparseTSF and FITS as baselines. In the FITS model, which employs a low-pass filter, we set the cutoff frequencies to the maximum and minimum harmonics provided in its source code, denoting them as FITS-T and FITS-L, respectively. Experiments on ETTh1, ETTm1, Electricity, and Weather datasets (Fig. 5) show that our model achieves the lowest MSE overall, with a parameter count falling between FITS-T and SparseTSF. While FITS-T reduces parameters by blocking more high-frequency information, it sacrifices performance. In contrast, our model not only uses fewer parameters than FITS-T but also matches or outperforms FITS-L. This demonstrates that our model design effectively balances performance and parameter efficiency for time series forecasting.
Table IV summarizes the parameter counts of leading models on the LTSF dataset, using an input look-back length of and an output length of . DiPE-Linear emerges as the most lightweight model, with parameter counts reduced by a factor of to compared to other baseline models across four LTSF datasets. This compact design not only significantly lowers computational requirements but also aligns with environmentally sustainable practices by reducing the carbon footprint.
| Model | ETTh1 | ETTm1 | Electricity | Weather |
| DiPE-Linear (Ours) | 4K | 4K | 17K | 16K |
| FITS | 154K | 60K | 411K | 572K |
| NLinear | 519K | 519K | 519K | 167M |
| RLinear | 519K | 519K | 519K | 167M |
| DLinear | 1M | 1M | 1M | 333M |
| iTransformer | 1M | 385K | 5M | 5M |
| TimesNet | 1M | 1M | 151M | 2M |
| PatchTST | 36M | 42M | 40M | 30M |
| Crossformer | 43M | 43M | 18M | 216K |
Fig. 2 further demonstrates the effectiveness of DiPE-Linear by comparing its average MSE and parameter counts with other baseline models across all prediction lengths and datasets. The results highlight that DiPE-Linear achieves state-of-the-art performance while maintaining a drastically smaller parameter count, underscoring its efficiency and practicality.
IV-B4 Effect of Training Data Size
To further investigate whether the model can perform effectively with reduced training data, we conducted experiments on the ETTh1 and ETTm1 datasets using different proportions of the training set. For the ETTh1 dataset, we compared our model with other FCs. Additionally, since the ETTm1 dataset contains more samples, we included nonlinear models for a broader comparison. As shown in Fig. 6, DiPE-Linear outperforms both FCs and nonlinear models across all proportions of the training set. Notably, on the ETTh1 dataset, our model surpasses the performance of other FC models trained on the full dataset while using only 40% of the training data.
IV-C Ablation Studies
IV-C1 Basic Components
To comprehensively validate the effectiveness of our model, we conducted ablation tests on two submodules: SFA and STA. The submodule IFM, which maps the input sequence space to the output sequence space, serves as the core component of our model and therefore cannot be ablated. Table V presents the results of the ablation experiments on the ETTh2 and Weather datasets with various prediction lengths. In both the ETTh2 and Weather datasets, the incorporation of the SFA and STA modules generally enhances the model’s prediction accuracy. Specifically, in the Weather dataset, the abundance of training samples ensures that each submodule effectively captures significant patterns, leading to a substantial improvement in prediction performance. This demonstrates the necessity of addressing the importance of both the frequency domain and time domain in our input data processing.
| Components | ETTh2 | Weather | ||||||||
| SFA | STA | IFM | 96 | 192 | 336 | 720 | 96 | 192 | 336 | 720 |
| ✗ | ✓ | ✓ | 0.284 | 0.331 | 0.348 | 0.382 | 0.146 | 0.191 | 0.243 | 0.315 |
| ✓ | ✗ | ✓ | 0.290 | 0.331 | 0.348 | 0.379 | 0.166 | 0.220 | 0.272 | 0.331 |
| ✗ | ✗ | ✓ | 0.295 | 0.340 | 0.352 | 0.387 | 0.175 | 0.228 | 0.274 | 0.332 |
| ✓ | ✓ | ✓ | 0.275 | 0.325 | 0.350 | 0.375 | 0.142 | 0.187 | 0.234 | 0.306 |
IV-C2 Low-rank weight sharing
In multivariate time series, our model must effectively capture the relationships between variables. We adopt a low-rank weight sharing approach, specifying the learning of sets of weights. Figure 7 presents the results of testing different values of on the Weather dataset with a prediction length of . When , our model is equivalent to weight sharing, meaning all variables use the same set of weights. When , our model learns a separate set of weights for each variable, treating the variables as independent. The experiments indicate that selecting an appropriate value for enables the model to achieve optimal performance.
We conducted an in-depth analysis to evaluate the necessity of low-rank weight sharing for each component of the model. Extensive experiments were performed on the ETTh2 and Weather datasets, covering multiple prediction horizons. For each component, we compared three configurations: shared weights, low-rank shared weights, and full-rank independent weights. As presented in Table VI, the low-rank weight sharing consistently delivered superior performance across all settings, highlighting its critical role in optimizing model efficacy.
| Components | ETTh2 | Weather | ||||||||
| SFA | STA | IFM | 96 | 192 | 336 | 720 | 96 | 192 | 336 | 720 |
| ✗ | ✓ | ✓ | 0.280 | 0.329 | 0.353 | 0.375 | 0.143 | 0.187 | 0.238 | 0.310 |
| ✓ | ✗ | ✓ | 0.279 | 0.326 | 0.354 | 0.374 | 0.142 | 0.187 | 0.238 | 0.309 |
| ✓ | ✓ | ✗ | 0.283 | 0.329 | 0.353 | 0.378 | 0.142 | 0.188 | 0.238 | 0.310 |
| F | F | F | 0.270 | 0.333 | 0.356 | 0.380 | 0.169 | 0.214 | 0.276 | 0.340 |
| ✓ | ✓ | ✓ | 0.275 | 0.325 | 0.350 | 0.375 | 0.142 | 0.187 | 0.234 | 0.306 |
IV-C3 SFALoss
The parameter in SFALoss governs the trade-off between the frequency domain loss and the time domain loss. To assess the influence of on model performance, we conducted experiments on the ETTh1 and ETTh2 datasets with a prediction length of . The results, illustrated in Figure 8, demonstrate the significant impact of on optimizing the model’s predictive accuracy across both datasets.
IV-D Interpretability: Case Studies
Benefiting from the disentangled representation of our model in both the frequency and time domains, our approach demonstrates superior interpretability. This interpretability facilitates user understanding of the model’s behavior, thereby enhancing trust in its predictive outcomes.
IV-D1 ETTh1
We select the ETTh1 dataset as a simple example. We choose an input length and a prediction horizon , ensuring the model has the minimum number of parameters. Additionally, within the ETTh1 dataset, we set the hyperparameter , which eliminates the router . Thus for each module, we learn only a single set of weights. The weight visualization is shown in Fig. 9. The SFA module explicitly preserves low-frequency information while emphasizing harmonic amplitudes in the high-frequency region. The STA module identifies the importance of each time point in the input sequence for future predictions. The IFM module learns distinct prediction patterns for harmonic frequencies (shown in red) and non-harmonic frequencies (shown in blue).
IV-D2 Weather
In the Weather dataset, different variables represent various meteorological data, such as temperature, pressure, humidity, wind direction, precipitation, etc.. We set as the hyperparameter, allowing the model to learn that the router can represent the correlations among different meteorological data variation patterns. The visualization is shown in Fig. 10.
We can further process the weights learned by the model to understand the knowledge it has captured. Taking the Router as an example, for each variable, the router weight can be viewed as the distribution of its predictive patterns across different weight groups. Consequently, we can compute the Jensen-Shannon Distance (JSD) between variables:
where are two variables, is a mixture distribution, and is the Kullback–Leibler (KL) divergence defined by 17:
Thus, we obtained the JSD Matrix based on the learnt router, as shown in Fig. 11.
IV-D3 Electricity
We further selected the Electricity dataset to comprehensively demonstrate the representations learned by DiPE-Linear. We chose an input length and a prediction horizon in the LSTF setting. In the Electricity dataset, the hyperparameter , meaning that four sets of weights are learned for each module. The parameter visualization is shown in Figure 10. Similar to the ETTh1 dataset, the SFA module learns to retain low-frequency information and high-frequency harmonics. The STA module clearly distinguishes the periodic variations between weekdays and weekends within a month. In the IFM module, four different frequency transformation patterns are also learned. Finally, the weight visualization of the Router illustrates the differences and similarities in predictive patterns across different variables in the dataset.
IV-E Complexity analysis
Our model retains the linear characteristics of FCs while achieving significant reductions in both parameter and computational complexity. In this section, we analyze the complexity of each module in our model in detail. For simplicity, we denote ambiguously both the input length and forecast horizon as .
IV-E1 Parameter Complexity
IV-E2 Computational Complexity
The computational cost of our model primarily arises from two components. First, the model’s parameter computations involve only element-wise multiplications, yielding a computational complexity of . Second, the rFFT and irFFT operations each have a computational complexity of . Therefore, the overall computational complexity of the model is .
A comparison of the parameter and computational complexities between our model and FCs (both with and without weight sharing) is provided in Table VII.
| DiPE-Linear (Ours) | FC (w/ sharing) | FC (w/o sharing) | |
| Param. | |||
| Compu. |
V Conclusion
This paper presents DiPE-Linear, a novel approach designed for LTSF in industrial applications. DiPE-Linear effectively addresses the challenges of Industry 5.0, including high-dimensional and high-resolution data as well as high-stakes scenarios, by leveraging disentangled, linear-complexity interpretable modules in both the time and frequency domains. The sophisticated design reduces parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear. Extensive experimental results demonstrate that DiPE-Linear consistently outperforms existing models, significantly enhancing scalability, computational efficiency, and predictive accuracy, particularly in complex industrial environments. In future work, we will enhance DiPE-Linear by incorporating interpretable nonlinear modules, while further exploring lightweight and efficient practical LSTF models.

### v2

@@
-Disentangled Interpretable Representation for Efficient Long-term Time Series Forecasting
+
 Abstract
-Industry 5.0 introduces new challenges for Long-term Time Series Forecasting (LTSF), characterized by high-dimensional, high-resolution data and high-stakes application scenarios.
-Against this backdrop, developing efficient and interpretable models for LTSF becomes a key challenge.
-Existing deep learning and linear models often suffer from excessive parameter complexity and lack intuitive interpretability.
-To address these issues, we propose DiPE-Linear, a Disentangled interpretable Parameter-Efficient Linear network.
-DiPE-Linear incorporates three temporal components: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM).
-These components alternate between learning in the frequency and time domains to achieve disentangled interpretability.
-The decomposed model structure reduces parameter complexity from quadratic in fully connected networks (FCs) to linear and computational complexity from quadratic to log-linear.
-Additionally, a Low-Rank Weight Sharing policy enhances the model’s ability to handle multivariate series.
-Despite operating within a subspace of FCs with limited expressive capacity, DiPE-Linear demonstrates comparable or superior performance to both FCs and nonlinear models across multiple open-source and real-world LTSF datasets, validating the effectiveness of its sophisticatedly designed structure.
-The combination of efficiency, accuracy, and interpretability makes DiPE-Linear a strong candidate for advancing LTSF in both research and real-world applications.
-The source code is available at https://github.com/wintertee/DiPE-Linear.
-Index Terms:
-Time series analysis, Time series forecasting, Efficient neural network, Interpretable artificial intelligenceI Introduction
-Long-term Time Series Forecasting (LTSF) is of significant importance in numerous domains, particularly in industrial applications such as cloud computing [1, 2], energy optimization [3, 4], and manufacturing [5, 6].
-Recent advancements in deep learning, including Recurrent Neural Networks (RNNs) [7], Convolutional Neural Networks (CNNs) [8, 9, 10, 11], Transformers [12, 13, 14, 15, 16, 17, 18, 19, 20], Multilayer Perceptrons (MLPs)[21, 22, 23, 24, 25, 26] and Fully Connected linear models (FCs) [27, 28, 29] have demonstrated promising performance for LTSF tasks.
-With the advent of Industry 5.0 and the era of big data, time-series data processing faces three primary challenges: high dimensionality, high resolution, and regulatory industrial scenarios.
-As regulatory emphasis on carbon footprint and trustworthiness increases[30], developing an efficient and interpretable time-series forecasting method has emerged as a critical challenge.
-Long-term forecasting for high-resolution time-series data remains challenging.
-Metrics like CPU load and network traffic in cloud environments are sampled at minute intervals, yet predictions must span days to support proactive resource allocation [31, 32, 20].
-Similarly, advanced sensors in industrial settings generate vast minute-level data, requiring accurate long-term forecasting for equipment monitoring and operational optimization [33].
-Existing methods face key limitations: RNNs struggle with parallelization [7], CNNs (e.g., TimesNet [11]) impose high computational costs, and Transformers, despite linear complexity improvements [16, 17], often overfit on long input sequences [27].
-Fully connected models offer simplicity and competitive performance but suffer from quadratic complexity and parameter redundancy (Figs.
-1(a), 1(b) and 1(c)).
-This raises the pivotal question: How can a model achieve parameter efficiency to effectively handle the challenges of long-term forecasting for high-resolution data?
-In high-stakes industrial scenarios, interpretability are indispensable requirements for achieving trustworthy AI [34, 35].
-In this context, domain experts must understand the model to trust its predictions.
-While leading deep learning models offer impressive precision, their black-box nature significantly undermines user trust.
-Fully connected linear models, by contrast, provide better interpretability by design.
-However, it remains challenging to fully and directly understand the features encoded by these models through their two-dimensional weight matrices, whether represented in the time domain (Figs.
-1(a) and 1(b)) or frequency domain (Fig.
-1(c)).
-This introduces the second key question: How can a model ensure interpretability, enabling domain experts to understand and trust its predictions in high-stakes industrial applications?
-To address these challenges, we revisit the performance of FCs in LTSF, seeking inductive biases to reduce complexity and enhance interpretability.
-Figs.
-1(a) and 1(b) reveals that highly correlated weights along the diagonal suggest parameter redundancy in the time domain, while the high noise in off-diagonal regions in the frequency domain as shown in Fig.
-1(c) indicates the independence of frequency components.
-This motivates our frequency independence hypothesis.
-Furthermore, time series often exhibit varying contributions of input time points to future predictions; for example, the most recent data typically plays a greater role in predicting future values.
-Thus, modeling the temporal significance of input sequences is crucial.
-Similarly, different frequency components contribute unequally[36], with significant frequencies often possessing higher signal-to-noise ratios and deserving more model attention.
-Finally, in multivariate time series, variable dependencies must be captured in an explainable manner, as variables may either share patterns or act independently.
-Based on the aforementioned motivations, we propose a novel Disentangled interpretable and Parameter-Efficient Linear model (DiPE-Linear) for long-term time series forecasting.
-As a deep linear model, it is structured into three interpretable linear sub-modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM).
-The SFA module extracts and enhances prominent frequencies in the frequency domain, and the STA module selects input time points in the time domain that are strongly correlated with the predicted sequence.
-Meanwhile, the IFM module assumes independence across different frequencies and directly maps the input series to the output series at each frequency component in frequency domain.
-By leveraging disentangled representations in both the time and frequency domains, the model achieves reduced complexity with linear parameter growth and log-linear computational scaling.
-Additionally, as shown in Fig.
-1(d), these representations enhance domain experts’ trust by providing clear insights into the learned features and the model’s reasoning process.
-These sub-modules employ a low-rank weight-sharing policy to efficiently model inter-variate relationships, reducing model complexity while maintaining strong performance.
-We further introduce a novel loss function, SFALoss, which is defined as a sum of time-domain loss and frequency-domain loss weighted by frequency-domain attention in the SFA module, guiding the model to prioritize significant frequencies with less noise.
-Additionally, by avoiding non-linearity within these modules, our model remains a linear model while achieving comparable or SOTA performance in LSTF with significantly fewer parameters, as shown in Fig.
-2.
-This lightweight design enhances model’s interpretability and establishes a more reliable and efficient prediction pipeline, offering valuable insights for future research.
+Long-term Time Series Forecasting (LTSF) is crucial across various domains, but complex deep models like Transformers are often prone to overfitting on extended sequences.
+Linear Fully Connected (FC) models have emerged as a powerful alternative, achieving competitive results with fewer parameters.
+However, their reliance on a single, monolithic weight matrix leads to quadratic parameter redundancy and an entanglement of temporal and frequential properties.
+To address this, we propose DiPE-Linear, a novel model that disentangles this monolithic mapping into a sequence of specialized, parameter-efficient modules.
+DiPE-Linear features three core components: Static Frequential Attention to prioritize critical frequencies, Static Time Attention to focus on key time steps, and Independent Frequential Mapping to independently process frequency components.
+A Low-rank Weight Sharing policy further enhances efficiency for multivariate data.
+This disentangled architecture collectively reduces parameter complexity from quadratic to linear and computational complexity to log-linear.
+Experiments on real-world datasets show that DiPE-Linear delivers state-of-the-art performance with significantly fewer parameters, establishing a new and highly efficient baseline for LTSF.
+Our code is available at https://github.com/wintertee/DiPE-Linear/
+1 Introduction
+Long-term Time Series Forecasting (LTSF) plays a crucial role in a wide array of fields, such as weather forecasting [1] and industrial manufacturing [27], making it a prominent focus of research in real-world applications.
+However, LTSF presents significant challenges due to the need for processing extended look-back window and forecasting horizon, which increases both the complexity of the forecasting task and the dimensionality of the model.
+Recent advancements in deep learning, such as Convolutional Neural Networks [12], and Transformers [20, 14, 11], have shown promise in tackling these challenges.
+However, as sequence length increases, the number of model parameters grows rapidly, significantly heightening the risk of overfitting [23].
+| (a) | DLinear [23] |
+| MSE=0.399 | |
+| #Params=18K |
+Trend
+Seasonal
+IR
+| (b) | FITS[21] |
+| MSE=0.385 | |
+| #Params=6.4K |
+Amplitude
+Phase
+IR
+| (c) | DiPE-Linear (Ours) |
+| MSE=0.384 | |
+| #Params=0.7K |
+SFA
+STA
+IFM
+IR
+In response to this overfitting challenge, Linear Fully Connected (FC) models have gained renewed attention.
+Recent research has shown that such FC models can match or even outperform complex nonlinear models, while requiring substantially fewer parameters [23, 7, 21, 10].
+However, despite addressing the parameter excess of nonlinear models, these FC models are often monolithic.
+They attempt to learn complex temporal dynamics using a single, dense weight matrix .
+This all-in-one design, which simultaneously captures frequential properties, temporal dependencies, and the input-output mapping, introduces its own critical issues: it not only leads to significant parameter redundancy, evidenced by its quadratic complexity, but also obscures the model’s inner working mechanisms.
+The inefficiency of this monolithic design can be observed directly.
+For instance, the weight matrix of a trained DLinear[23] model (Fig.
+1a) exhibits distinct horizontal and diagonal stripes.
+These stripes represent the same pattern repeating at different periodic lags, revealing a massive parameter redundancy—the model is inefficiently storing the same temporal dependency multiple times.
+Meanwhile, the FITS[21] off-diagonal amplitude and phase visualization (Fig.
+1b) appears noisy.
+These visualizations suggest that while FC models attempt to learn structural properties, they do so implicitly and inefficiently.
+We identify this core problem as entanglement: the properties are wastefully embedded within a single, large, dense matrix.
+This motivates our core idea: to replace the monolithic matrix with an explicit factorization, or disentanglement, into a sequence of specialized, interpretable, and parameter-efficient linear modules.
+We disentangle the core temporal dynamics into three key priors: (1) Frequential Filtering: In recognition of the fact that time series data are typically decomposed into trend and seasonal components [20], we incorporate the Static Frequential Attention (SFA) module to prioritize predictable frequencies.
+(2) Temporal Importance: Not all historical time steps are equally important.
+This motivates the Static Time Attention (STA) module to focus on pivotal time steps.
+(3) Independent Mapping: To model these periodicities holistically and efficiently through a time-invariant mapping (thus avoiding the explicit redundancy seen in Fig.
+1), we propose the Independent Frequential Mapping (IFM) approach.
+Furthermore, to efficiently capture inter-variable dependencies [15] in the multivariate setting, we introduce a (4) Low-rank Weight Sharing policy.
+Our explicit factorization (Fig.
+1c) achieves a cleaner, more efficient representation, directly addressing the redundancy observed in Fig.
+1a and 1b.
+Building upon the above-mentioned motivation, we propose a novel Disentangled Parameter-Efficient Linear model (DiPE-Linear) for LTSF.
+In this work, "disentangled" refers to the model’s architecture which factorizes the complex, monolithic FC weight matrix into a sequence of specialized, interpretable, and parameter-efficient linear modules (SFA, STA, and IFM).
+Each module addresses a distinct aspect of temporal modeling (frequential filtering, temporal importance weighting, and frequential mapping, respectively).
+These linear sub-modules collectively reduce the FC’s quadratic parameter complexity to linear complexity and quadratic computational complexity to log-linear.
+This lightweight, fully linear design achieves state-of-the-art performance with minimal parameters, as shown in Fig.
+1 and 2, establishing an efficient and reliable pipeline for time series forecasting.
 In summary, our contributions are multifold:
-We propose DiPE-Linear, a deep linear model that achieves efficiency, significantly reducing parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear.
-Disentangled representation of model components in both time and frequency domain offers superior interpretability, thereby enhancing its trustworthiness for industrial applications.
-Extensive experiments demonstrate that our method outperforms existing approaches, offering superior performance, efficiency, and interpretability.
-II Related Work
-II-A Efficient Architectures for Long-Term Forecasting
-Long-term time series forecasting (LTSF) requires models capable of efficiently handling long sequences.
-Early Transformer-based models [12, 13, 14, 15, 16, 17] progressively reduced computational complexity from to .
-However, due to the inherently complex structure of Transformers, these models still contain a large number of parameters and are highly prone to overfitting when dealing with excessively long input sequences in LTSF tasks[27].
-DLinear[27] was the first to propose using a simple single-layer fully connected network to replace the complex Transformer structure.
-Although the fully connected layer has a computational complexity of , its simple structure results in significantly lower complexity compared to other Transformer-based models.
-Nevertheless, DLinear still suffers from notable parameter redundancy.
-FITS[29] addressed this by employing low-pass filters to discard high-frequency information in the frequency domain, drastically reducing the number of parameters in the fully connected network to the scale of 10k.
-SparseTSF[37], on the other hand, introduced Cross-Period Sparse Forecasting, which achieves a parameter count on the scale of 1k through downsampling and cross-period parameter sharing.
-However, both FITS and SparseTSF sacrifice accuracy to achieve parameter reduction[38, 37].
-Further exploration into addressing parameter redundancy in fully connected models remains an open research question.
-II-B Exploiting Frequency-Domain Characteristics
-Recent advancements in time series forecasting have increasingly leveraged frequency-domain knowledge to improve performance.
-One prominent approach incorporates frequency-domain information as auxiliary features to enhance the modeling of long-term dependencies.
-For example, FEDformer[17] introduces a DFT-based frequency-enhanced attention mechanism, where attention weights are computed from the spectra of queries and keys, with the weighted sum calculated directly in the frequency domain.
-Similarly, FiLM[39] applies Fourier analysis to retain historical information while filtering out noise.
-Alternatively, some methods operate entirely in the frequency domain.
-FreTS[40] employs a frequency-domain MLP to capture both channel-wise and temporal dependencies, while FITS[29] adopts a fully connected network specifically designed for spectral representations.
-Another promising direction incorporates frequency-domain loss functions into time-domain training objectives to capture spectral patterns more effectively.
-For example, FTMixer [26] and FreDF [41] add frequency-domain loss components on top of time-domain losses, enhancing the model’s ability to align with the ground truth in the frequency domain.
-II-C Channel-Independence for Multivariate Time Series
-DLinear[27] was the first to introduce the Channel-Independent strategy for multivariate time series forecasting (the terms ”channel” and ”variate” are used interchangeably throughout this paper for simplicity).
-Although this approach ignores causal relationships between variables, it achieves significant performance improvements.
-PatchTST[19] further validated this idea in Transformer-based models, demonstrating that independently processing variables while sharing the same model weights across variables can be highly effective.
-However, RLinear[28] shows that for variables with distinct periodic characteristics, using independent weights for each variable can further enhance performance.
-Subsequent research has focused on finding an optimal balance between weight sharing and weight independence.
-For instance, MLinear[42] dynamically adjusts between Weight Sharing and Weight Independent strategies based on the temporal semantics of different sequences, while CSC[43] reassigns weight-variable mappings during the validation phase to achieve variable clustering.
-Nevertheless, these methods come with limitations.
-MLinear requires calculating both Weight Sharing and Weight Independent outputs before combining them, which increases computational overhead.
-CSC, on the other hand, relies on a strong prior assumption that variables clustered together should share the same weights, which may not hold universally.
-III Method
-DiPE-Linear is constructed from several essential components.
-We begin by outlining the preliminaries related to TSF and the Fourier Transform, providing the foundation for our approach.
-Next, we present the architecture of DiPE-Linear, detailing its three core modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM), under assumption that the input series is univariate.
-Then, we describe how the Low-rank Weight Sharing method generalizes those modules to multivariate time series.
-Lastly, we introduce SFALoss, a novel loss function that combines weighted losses in both the frequency and time domains, guiding the model to prioritize informative patterns while enhancing predictive performance.
-III-A Preliminaries
-III-A1 Problem Definition
-The multivariate TSF problem can be formally defined as follows: Given a historical time series , where is the number of variates and is the look-back length of input time series, the objective is to predict the future time series , where is the forecasting horizon.
-III-A2 Discrete Fourier Transform
-Time series data can be viewed as discrete samples of a continuous signal over time.
-In many signal processing and time-series forecasting applications, the Discrete Fourier Transform (DFT) is a commonly used method to analyze the frequency components of such discrete signals.
-For a discrete-time sequence of length , the DFT transforms the time-domain signal into its frequency-domain representation.
-This transformation is particularly useful for identifying underlying periodic patterns or extracting frequency components that contribute to the signal.
-The DFT of a time series is defined as:
-Once the signal is represented in the frequency domain, various operations, such as filtering or modifying specific frequency components, can be applied.
-The inverse Discrete Fourier Transform (iDFT) allows for transforming the modified frequency-domain signal back to the time domain.
-The iDFT is expressed as:
-This transformation provides the flexibility to manipulate signals in the frequency domain and revert them back to the time domain for further analysis or application.
-The ability to move between domains seamlessly is essential for efficient signal processing.
-III-A3 Duality Between Time and Frequency Domains
-The DFT exhibits a fundamental property known as duality between the time and frequency domains.
-This duality manifests as convolution in the time domain being equivalent to multiplication in the frequency domain, and vice versa.
-For two time-domain signals, and , their convolution:
-has a frequency-domain equivalent:
-Pointwise multiplication in the time domain, on the other hand, corresponds to convolution in the frequency domain.
-This relationship simplifies many operations, particularly in tasks where convolution is computationally intensive in the time domain.
-Thus, the DFT provides a powerful tool for analyzing time series in the frequency domain, leveraging the duality between time and frequency domains to simplify complex operations such as convolution.
-The ability to transition between domains efficiently enables advanced signal processing techniques for both single-channel and multichannel data.
-III-B Static Frequential Attention
-Extracting frequencies is a common practice, as previous FITS [29] employed a low-pass filter to eliminate high-frequency signals, while FedFormer [17] randomly selects a subset of frequencies as key features.
-However, such strong prior assumptions may lead the model to discard useful information.
-Therefore, we aim for the model to learn which frequency components to focus on.
-Specifically, to ensure the model’s capability to select and amplify significant frequencies, we introduce SFA module as a frequency-domain filter.
+We propose DiPE-Linear, a novel disentangled parameter-efficient linear model for LTSF.
+Its innovative design, motivated by an analysis of entanglement in existing FCs, features specialized sub-modules that improve performance and offer clear interpretability.
+The proposed DiPE-Linear significantly reduces the parameter complexity (from quadratic to linear) as well as computational complexity (to log-linear) without sacrificing accuracy.
+Extensive experiments demonstrate that DiPE-Linear achieves performance comparable to or surpassing other state-of-the-art models, validating its effectiveness and efficiency.
+2 Related Work
+2.0.1 Efficient Long-Term Forecasting.
+LTSF demands models that handle long sequences efficiently.
+Early Transformer-based models [5, 4, 24, 20, 26] reduced computational complexity from to , but their complex structure leads to a large number of parameters, causing overfitting on long inputs [23].
+DLinear [23] proposed a simpler, single-layer fully connected network with complexity, offering lower overall complexity than Transformer-based models, though it still suffers from parameter redundancy.
+FITS [21] reduced parameter count by applying low-pass filters in the frequency domain, while SparseTSF [10] introduced Cross-Period Sparse Forecasting to achieve further parameter reduction.
+However, both methods trade off accuracy to reduce parameters [16, 10], and addressing parameter redundancy in FC models remains an open challenge.
+2.0.2 Exploiting Frequency-Domain Characteristics.
+Recent advances in time series forecasting have leveraged frequency-domain techniques to enhance performance.
+FEDformer [26] introduced a DFT-based frequency-enhanced attention mechanism, while FiLM [25] used Fourier analysis to retain historical information and reduce noise.
+Other methods, such as FreTS [22] and FITS [21], work entirely in the frequency domain, with FreTS using a frequency-domain MLP for both temporal and channel-wise dependencies.
+Additionally, frequency-domain loss functions have been integrated into time-domain objectives.
+For example, FTMixer [8] and FreDF [17] use frequency-domain loss components to improve alignment with the ground truth.
+2.0.3 Channel Dependence for Multivariate Forecasting.
+DLinear [23] introduced the Channel-Independent strategy for multivariate time series forecasting, which, despite ignoring causal relationships, led to notable performance gains.
+PatchTST [14] further validated this approach in Transformer-based models, showing that independently processing variables while sharing model weights across them can be effective.
+RLinear [7] demonstrated that using independent weights for variables with distinct periodicities can further improve performance.
+Subsequent research has sought an optimal balance between weight sharing and independence.
+MLinear [6] dynamically adjusts strategies based on temporal semantics, while CSC [15] reassigns weight-variable mappings during validation to cluster variables.
+However, MLinear introduces extra computational overhead, and CSC relies on assumptions about weight sharing that may not always hold.
+3 Method
+We first define the multivariate LTSF problem, then describe the DiPE-Linear architecture, focusing on its three core modules: Static Frequential Attention (SFA), Static Temporal Attention (STA), and Independent Frequential Mapping (IFM), assuming univariate time series.
+We then explain how the Low-rank Weight Sharing method extends these modules to multivariate time series.
+Finally, we introduce SFALoss, a loss function that combines weighted losses in both frequency and time domains to prioritize predictable frequencies and improve model performance.
+3.1 Problem Definition
+The multivariate LTSF problem is formally defined as follows: Given a historical time series , where denotes the number of variates and represents the look-back length, the goal is to predict the future time series , where is the forecasting horizon.
+3.2 Static Frequential Attention
+We introduce the SFA module as a frequency-domain filter to enable the model to selectively enhance and amplify relevant frequencies.
 Additionally, to preserve the temporal structure of the series and avoid interference with subsequent temporal feature extraction, we constrained this filter to be a zero-phase filter.
@@
 where represents the input univariate series , denotes the rFFT, is the irFFT, is the learnable static frequential attention map, and denotes element-wise multiplication.
-III-C Static Temporal Attention
-Similarly, to enable the model to have the capability of capturing important input time points in the temporal domain, we introduce the STA as a second component of our model.
+3.3 Static Temporal Attention
+We apply SFA before STA based on the hypothesis that the model should first isolate the most predictable, core frequencies of the signal (SFA), and then identify the critical time steps within that filtered signal (STA).
+Reversing the order might cause the STA module to focus on temporally salient but high-frequency noise.
+Following this, to enable the model to have the capability of capturing important input time points in the temporal domain, we introduce the STA as a second component of our model.
 This module enables the input time series to undergo element-wise multiplication with a learned temporal attention map, effectively assigning appropriate importance to relevant historical time points and thereby enhancing the model’s ability to capture temporal dependencies.
@@
 where is the learnable time-domain attention map.
-III-D Independent Frequential Mapping
-While both the SFA and STA modules function as part of the input series pre-processing, the IFM module directly maps the input historical series to the forecasted output.
-To process the time series as a whole, leveraging the spectrum is a natural approach, as it encapsulates all the information about the entire sequence and enables better handling of periodic components.
-Properly and efficiently processing the spectrum remains a key challenge.
-Fig.
-1(c) illustrates that the two-dimensional matrix used by FITS [29] to handle the spectrum is sparse, with only the diagonal weights being significant.
-This suggests that, for a time series, there is relatively strong independence between different frequencies, allowing each frequency to be handled separately.
-We thus treat each frequency independently for two key reasons: (1) Apart from harmonic frequencies, there is generally little correlation observed between different frequencies in time series data.[41] (2) Although harmonic frequencies tend to exhibit strong correlations, we avoid explicitly modeling these inter-frequency dependencies to mitigate potential multicollinearity in linear regression.
-Therefore, for each frequency component, we apply a complex-valued multiply-accumulate operation independently.
+3.4 Independent Frequential Mapping
+While the SFA and STA modules are part of the input series pre-processing, the IFM module directly maps the input historical series to the forecasted output under the assumption of independence between different frequencies.
+We acknowledge that this assumption is a theoretical simplification, as real-world signals often exhibit complex inter-frequency dependencies (e.g., harmonics [17]).
+However, we adopt this assumption as a deliberate design choice.
+Explicitly modeling these intricate dependencies within a linear framework could increase the risk of multicollinearity and overfitting.
+Therefore, our approach acts as an effective inductive bias.
+By enforcing frequency-wise independence, the IFM module learns a 1D convolution (Eq.
+6) that is inherently time-invariant and phase-preserving (relative to each frequency component), which are desirable properties for modeling stationary seasonal patterns.
+This compels the model to focus on the primary predictive information within each frequency component rather than capturing potentially spurious cross-frequency correlations.
+As our extensive experiments on diverse datasets (see Section 4) demonstrate, this design choice consistently leads to state-of-the-art performance, suggesting it enhances model generalization and robustness rather than limiting its accuracy.
+Following this assumption, for each frequency component, we apply a complex-valued multiply-accumulate operation independently.
 To transform the input space to the output space , we apply zero-padding to the input series, extending its length to before performing the rFFT.
@@
 The forecasted output is thus defined as:
-where is the complex-valued weight and is the complex-valued bias.
-This process can be equivalently described as a 1D convolution in time domain, as explained in equation 3 and 4:
+where is the complex-valued weight of the IFM module, and is the complex-valued bias.
+According to the Convolution Theorem, this process can be equivalently represented as a 1D convolution in the time domain:
 where is the convolution kernel in time domain.
@@
 This larger kernel allows each output to capture a global receptive field, enabling the model to incorporate information from the entire input sequence.
-III-E Low-rank Weight Sharing
-Weight sharing is a common practice in time series forecasting models [27, 29, 19], based on the assumption that different channels of a time series exhibit similar patterns.
-However, this assumption often does not hold.
-For example, in Weather datasets, variables such as temperature, pressure, and rainfall may follow distinct patterns.
-In such cases, training weights independently for each channel can improve performance [28], but this approach significantly increases the number of parameters by a factor of , making the model complexity significantly higher.
-To address this, we propose a more efficient alternative.
-By clustering variables based on their pattern similarities, we can reduce the number of weight sets, balancing the trade-off between model efficiency and performance.
-This clustering approach not only enhances computational efficiency by reducing the number of learned parameters but also improves accuracy, as the shared weights capture more robust patterns.
-Specifically, well-structured clustering ensures that the learned representations generalize better across similar variables.
-Inspired by mixture-of-experts[44] and dynamic convolution[45], we introduce a novel Low-rank Weight Sharing architecture, as shown in Fig.
-4, leveraging these clustered patterns to optimize both parameter efficiency and predictive performance.
+3.5 Low-rank Weight Sharing
+Weight sharing is commonly used in time series forecasting models [23, 21, 14] under the assumption that different channels exhibit similar patterns.
+However, this assumption does not hold in datasets like Weather, where variables such as temperature, pressure, and rainfall follow distinct patterns.
+While training weights independently for each channel can improve performance [7], it significantly increases model complexity.
+To address this, we propose selecting a lower rank based on pattern similarities, reducing the number of weight sets and balancing efficiency with performance.
+Inspired by mixture-of-experts [13] and dynamic convolution [2], we introduce a novel Low-rank Weight Sharing architecture, optimizing parameter efficiency and predictive performance.
 Specifically, for a multivariate time series with variables, we introduce a hyperparameter , where denotes the number of independent weight sets to be learned, with .
-For each component, distinct sets of weights are learned, alongside a static routing matrix.
-The low-rank routing matrix is designed to linearly combine the weights and assign each variable to an appropriate weight set.
-We regularize the static routing matrix using the Softmax function, where is the Softmax temperature:
+For each component, distinct sets of weights are learned, alongside a learnable routing matrix.
+The low-rank routing matrix is a model parameter designed to linearly combine the weights and assign each variable to an appropriate weight set.
+We regularize this routing matrix using the Softmax function, where is the Softmax temperature:
 We denote as the set of mappings representing any given module, corresponding to the independent weight sets learned. For variate , the weight actually applied to channel is:
-III-F SFALoss
-Building upon the insights of FreDF [41], we introduce SFALoss (denoted as ), which combines Weighted Mean Absolute Error (WMAE) in the frequency domain () and MSE in the time domain ().
-For the frequency domain component (), our SFA module selectively subtracts key frequencies while attenuating noisy frequencies.
-To refine this process, we modulate the frequency-domain loss through an element-wise multiplication with the SFA weighting factor, , effectively filtering out loss contributions from less relevant frequencies.
-The WMAE loss in the frequency domain is formulated as follows:
+3.6 SFALoss
+Building on the insights from FreDF [17], we introduce SFALoss (denoted as ), which integrates the Weighted Mean Absolute Error (WMAE) in the frequency domain () and Mean Squared Error (MSE) in the time domain ().
+We empirically found this combination to be most effective: WMAE for provides a more robust optimization signal that is less sensitive to large magnitude errors in a few dominant frequencies, while the standard MSE for ensures overall predictive accuracy in the time domain.
+The frequency-domain loss () is modulated by element-wise multiplication with the SFA weighting factor, , allowing prioritizing relevant frequencies:
 where and are respectively ground truth and predicted future time series in frequency-domain.
 While computing the loss, is detached from the computational graph, ensuring that no gradients are backpropagated through it.
+This crucial step prevents a trivial solution where the SFA module might learn to suppress (i.e., set weights to zero for) frequencies that are difficult to predict, rather than learning the frequencies most relevant to the primary time-domain forecasting task.
 We also preserve the MSE loss in time domain :
@@
 where is hyperparameter that balances the contribution of the two loss terms, allowing us to achieve a trade-off between them.
-IV Experiment
-IV-A Experimental Setup
-IV-A1 Dataset
-To comprehensively evaluate the performance and generalization capability of our model, we selected a diverse set of eight datasets, including six for LTSF and two for short-term time series forecasting (STSF).
-These datasets come from various domains, with different sizes and sampling frequencies, offering a robust benchmark for assessing the versatility and scalability of our approach across a wide range of forecasting tasks.
-Details are shown in Table I.
-For the LTSF task, we utilized six public datasets:
-The ETT111https://github.com/zhouhaoyi/ETDataset dataset, including ETTh1, ETTh2, ETTm1 and ETTm2, comprises load and oil temperature data from two power transformers at different stations.
-The Electricity222https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 dataset includes electricity consumption data, measured in kilowatt-hours (kWh), from 321 clients, spanning from 2012 to 2014.
-The Weather333https://www.bgc-jena.mpg.de/wetter/ dataset consists of 21 meteorological indicators, sampled every 10 minutes, covering the entire year of 2020.
-Moreover, to validate the effectiveness of the model in real-world production environments, we selected two workload dataset collected from ByteDance’s cloud servers, serveing as private LTSF datasets:
-The FaaS includes 15-day query-per-second (QPS) data from the public cloud platform.
-The IaaS includes 45-day CPU usage data from the private cloud platform.
-To further demonstrate the generalizability of our model, we extended our evaluation to two short-term forecasting tasks using the following datasets:
-The Illness444https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html dataset records weekly influenza-like illness cases as reported by the Centers for Disease Control and Prevention (CDC) in the United States, from 2002 to 2021.
-The M5555https://www.kaggle.com/competitions/m5-forecasting-accuracy/ dataset contains hierarchical sales data from Walmart, covering stores in three U.S. states.
-| Category | Dataset | Domain | Variates | Samples | Interval |
-| LTSF | ETTh1 | Industry | 7 | 17420 | 1 hour |
-| ETTh2 | Industry | 7 | 17420 | 1 hour | |
-| ETTm1 | Industry | 7 | 69680 | 15 min | |
-| ETTm2 | Industry | 7 | 69680 | 15 min | |
-| Electricity | Energy | 321 | 26304 | 1 hour | |
-| Weather | Environment | 21 | 52696 | 1 hour | |
-| FaaS | Cloud Computing | 1014 | 23041 | 1 min | |
-| IaaS | Cloud Computing | 168 | 69045 | 1 min | |
-| STSF | Illness | Medical | 7 | 966 | 1 week |
-| M5 | Market | 30 | 1941 | 1 day |
-| DiPE-Linear | FITS | RLinear | DLinear | Nlinear | |||||||
-| (Ours) | (ICLR-24) | (arXiv-23) | (AAAI-23) | (AAAI-23) | |||||||
+4 Experiment
+4.1 Experimental Setup
+4.1.1 Datasets.
+We evaluate our model on 10 real-world datasets, comprising 8 for Long-term (LTSF) and 2 for Short-term (STSF) forecasting.
+The LTSF datasets include ETT (ETTh1, ETTh2, ETTm1, ETTm2) [24], Weather [24], Electricity, and two private cloud service datasets (IaaS, FaaS).
+The STSF datasets are M5 and Illness [20].
+Further preprocessing details are available in the accompanying code.
+4.1.2 Implementation Details.
+Models are trained for 50 epochs with a learning rate.
+Batch sizes are set to 64 (LTSF) and 8 (STSF).
+Forecasting horizons are for LTSF and for STSF.
+Input sequence lengths are 720 (public LTSF), 2880 (private LTSF), and 60 (STSF).
+We set for Electricity and Weather, otherwise, and apply linear annealing for the softmax temperature [2].
+4.1.3 Baselines and Metrics.
+We benchmark our approach against leading linear models, including FITS [21], RLinear [7], DLinear, and NLinear [23].
+We also compare with established non-linear models (e.g., iTransformer [11], PatchTST [3], TimesNet [19]) and lightweight baselines (SparseTSF [10], FITS [21]).
+Performance is evaluated using MSE and MAE.
+4.2 Main Results
+4.2.1 Versus FC Models.
+| DiPE-Linear (Ours) | FITS (2024) | RLinear (2023) | DLinear (2023) | Nlinear (2023) | |||||||
 | Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
@@
 | Best count | 32 | 26 | 7 | 5 | 7 | 7 | 6 | 1 | 9 | 6 | |
-| DiPE-Linear | iTransformer | PatchTST† | TimesNet† | Crossformer† | MICN† | FEDformer† | |||||||||
-| (Ours) | (ICLR-24) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICLR-23) | (ICML-22) | |||||||||
-| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
-| ETTh1 | 96 | 0.369 | 0.393 | 0.386 | 0.405 | 0.370 | 0.399 | 0.384 | 0.402 | 0.386 | 0.429 | 0.396 | 0.427 | 0.376 | 0.415 |
-| 192 | 0.407 | 0.415 | 0.424 | 0.440 | 0.413 | 0.421 | 0.557 | 0.436 | 0.419 | 0.444 | 0.430 | 0.453 | 0.423 | 0.446 | |
-| 336 | 0.424 | 0.427 | 0.449 | 0.460 | 0.422 | 0.436 | 0.491 | 0.469 | 0.440 | 0.461 | 0.433 | 0.458 | 0.444 | 0.462 | |
-| 720 | 0.409 | 0.439 | 0.495 | 0.487 | 0.447 | 0.466 | 0.521 | 0.500 | 0.519 | 0.524 | 0.474 | 0.508 | 0.469 | 0.492 | |
-| ETTh2 | 96 | 0.275 | 0.336 | 0.297 | 0.348 | 0.274 | 0.336 | 0.340 | 0.374 | 0.276 | 0.338 | 0.289 | 0.357 | 0.332 | 0.374 |
-| 192 | 0.325 | 0.372 | 0.371 | 0.403 | 0.339 | 0.379 | 0.402 | 0.414 | 0.345 | 0.382 | 0.409 | 0.438 | 0.407 | 0.446 | |
-| 336 | 0.350 | 0.393 | 0.404 | 0.428 | 0.329 | 0.380 | 0.452 | 0.452 | 0.375 | 0.411 | 0.417 | 0.452 | 0.400 | 0.447 | |
-| 720 | 0.375 | 0.415 | 0.424 | 0.444 | 0.379 | 0.422 | 0.462 | 0.468 | 0.408 | 0.446 | 0.426 | 0.473 | 0.412 | 0.469 | |
-| ETTm1 | 96 | 0.309 | 0.350 | 0.300 | 0.353 | 0.290 | 0.342 | 0.338 | 0.375 | 0.316 | 0.373 | 0.314 | 0.360 | 0.326 | 0.390 |
-| 192 | 0.339 | 0.369 | 0.345 | 0.382 | 0.332 | 0.369 | 0.371 | 0.387 | 0.377 | 0.411 | 0.359 | 0.387 | 0.365 | 0.415 | |
-| 336 | 0.367 | 0.386 | 0.374 | 0.398 | 0.369 | 0.392 | 0.410 | 0.411 | 0.431 | 0.442 | 0.398 | 0.413 | 0.392 | 0.425 | |
-| 720 | 0.416 | 0.413 | 0.429 | 0.430 | 0.416 | 0.420 | 0.478 | 0.450 | 0.600 | 0.547 | 0.459 | 0.464 | 0.446 | 0.458 | |
-| ETTm2 | 96 | 0.162 | 0.252 | 0.175 | 0.266 | 0.165 | 0.255 | 0.187 | 0.267 | 0.421 | 0.461 | 0.178 | 0.273 | 0.180 | 0.271 |
-| 192 | 0.216 | 0.289 | 0.242 | 0.312 | 0.220 | 0.292 | 0.249 | 0.309 | 0.503 | 0.519 | 0.245 | 0.316 | 0.252 | 0.318 | |
-| 336 | 0.268 | 0.324 | 0.282 | 0.340 | 0.274 | 0.329 | 0.321 | 0.351 | 0.611 | 0.580 | 0.295 | 0.350 | 0.324 | 0.364 | |
-| 720 | 0.353 | 0.379 | 0.378 | 0.398 | 0.362 | 0.385 | 0.497 | 0.403 | 0.996 | 0.750 | 0.389 | 0.406 | 0.410 | 0.420 | |
-| ECL | 96 | 0.132 | 0.228 | 0.132 | 0.229 | 0.129 | 0.222 | 0.168 | 0.272 | 0.187 | 0.283 | 0.159 | 0.267 | 0.186 | 0.302 |
-| 192 | 0.148 | 0.245 | 0.151 | 0.246 | 0.147 | 0.240 | 0.184 | 0.289 | 0.258 | 0.330 | 0.168 | 0.279 | 0.197 | 0.311 | |
-| 336 | 0.162 | 0.261 | 0.167 | 0.264 | 0.163 | 0.259 | 0.198 | 0.300 | 0.323 | 0.369 | 0.196 | 0.308 | 0.213 | 0.328 | |
-| 720 | 0.198 | 0.296 | 0.194 | 0.286 | 0.197 | 0.290 | 0.220 | 0.320 | 0.404 | 0.423 | 0.203 | 0.312 | 0.233 | 0.344 | |
-| Weather | 96 | 0.142 | 0.201 | 0.159 | 0.208 | 0.149 | 0.198 | 0.172 | 0.220 | 0.153 | 0.217 | 0.161 | 0.226 | 0.238 | 0.314 |
-| 192 | 0.187 | 0.253 | 0.200 | 0.248 | 0.194 | 0.241 | 0.219 | 0.261 | 0.197 | 0.269 | 0.220 | 0.283 | 0.275 | 0.329 | |
-| 336 | 0.234 | 0.293 | 0.253 | 0.289 | 0.245 | 0.282 | 0.280 | 0.306 | 0.252 | 0.311 | 0.275 | 0.328 | 0.339 | 0.377 | |
-| 720 | 0.306 | 0.348 | 0.321 | 0.338 | 0.314 | 0.334 | 0.365 | 0.359 | 0.318 | 0.363 | 0.311 | 0.356 | 0.389 | 0.409 | |
-| Best count | 16 | 14 | 1 | 1 | 8 | 11 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
-IV-A2 Implementation details
-All experiments are implemented in PyTorch [46] on an NVIDIA 4090D GPU.
-The models are trained using the Adam optimizer [47] with a learning rate of for epochs.
-Batch sizes are set to 64 and 8 for LTSF and STSF datasets, respectively.
-For Electricity and Weather datasets, , and otherwise.
-Input sequence lengths are , , and for public LTSF, private LTSF, and STSF datasets, respectively.
-Following [45], we apply linear annealing for the softmax temperature in Low-rank Weight Sharing to ensure balanced learning during early epochs.
-IV-A3 Baselines
-In this work, we focus on evaluating and comparing linear models, benchmarking our approach against leading methods such as FITS [29], RLinear [28], DLinear, and NLinear [27].
-To further demonstrate the effectiveness of our model, we also compare it with nonlinear deep learning models, including iTransformer [20], PatchTST [48], TimesNet [11], Crossformer [18], MICN [10], and FEDformer [17], showing that our model achieves state-of-the-art or comparable performance across both linear and nonlinear baselines.
-Additionally, SparseTSF [37] and FITS [29] are selected as parameter-efficient baselines to highlight the efficiency of our model design.
-IV-A4 Evaluation Metrics
-We follow the established practices in prior works to evaluate forecasting performance using two primary metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE).
-IV-B Main Results
-IV-B1 Versus FCs
-We first compared our model DiPE-Linear with other leading FCs.
-To ensure a fair evaluation, all models were trained using the same trainer settings.
-Throughout the training process, we carefully monitored each model to ensure convergence under these consistent conditions.
-All models use the optimal model hyperparameters provided in the authors’ original implementations.
-Table II presents the evaluation results for the linear models on the LTSF and STSF datasets.
+We first compared our model DiPE-Linear with other leading FC models.
+To ensure a fair evaluation, all models were trained to converge under consistent training settings and using optimal hyperparameters from the authors’ original implementations.
+Table 1 presents the evaluation results for FC models on the LTSF and STSF datasets.
 Across all eight datasets, DiPE-Linear consistently delivers comparable or even superior performance.
@@
 On larger datasets (ETTm1, ETTm2, Electricity, and Weather), DiPE-Linear achieves results comparable to other linear models.
-One notable advantage of DiPE-Linear is its substantially reduced parameter count compared to FCs.
-By operating within a low-dimensional manifold in the FCs’ model space, DiPE-Linear achieves performance comparable to or better than FCs, suggesting that our carefully designed model space contains the optimal solution to TSF tasks.
-This highlights the robustness and effectiveness of our architecture in representing the essential solution space of FCs with minimal complexity.
-IV-B2 Versus nonlinear models
-Next, we compare DiPE-Linear with several nonlinear models, including Transformer-based models amd CNN-based models.
-Table III presents the evaluation results for these nonlinear models on public LTSF datasets.
-For private LTSF datasets, excessively long input lengths result in high memory usage for nonlinear models, leading to unacceptably high deployment costs in production environments.
-For STSF datasets, the limited amount of training data makes the models highly prone to overfitting, making training challenging.
-Therefore, we do not report model performance on these two types of datasets.
-To ensure a fair comparison, all nonlinear models were trained with input lengths , and the best results were selected for evaluation.
-Our model demonstrates a slight performance advantage over PatchTST, while requiring only times the parameter count.
-Furthermore, our model significantly outperforms other nonlinear models, highlighting its competitive advantages in both efficiency and performance.
-IV-B3 Parameter complexity and efficiency
-Recent studies have focused on parameter-efficient models, and we selected SparseTSF and FITS as baselines.
-In the FITS model, which employs a low-pass filter, we set the cutoff frequencies to the maximum and minimum harmonics provided in its source code, denoting them as FITS-T and FITS-L, respectively.
-Experiments on ETTh1, ETTm1, Electricity, and Weather datasets (Fig.
-5) show that our model achieves the lowest MSE overall, with a parameter count falling between FITS-T and SparseTSF.
-While FITS-T reduces parameters by blocking more high-frequency information, it sacrifices performance.
-In contrast, our model not only uses fewer parameters than FITS-T but also matches or outperforms FITS-L.
-This demonstrates that our model design effectively balances performance and parameter efficiency for time series forecasting.
-Table IV summarizes the parameter counts of leading models on the LTSF dataset, using an input look-back length of and an output length of .
-DiPE-Linear emerges as the most lightweight model, with parameter counts reduced by a factor of to compared to other baseline models across four LTSF datasets.
-This compact design not only significantly lowers computational requirements but also aligns with environmentally sustainable practices by reducing the carbon footprint.
-| Model | ETTh1 | ETTm1 | Electricity | Weather |
-| DiPE-Linear (Ours) | 4K | 4K | 17K | 16K |
-| FITS | 154K | 60K | 411K | 572K |
-| NLinear | 519K | 519K | 519K | 167M |
-| RLinear | 519K | 519K | 519K | 167M |
-| DLinear | 1M | 1M | 1M | 333M |
-| iTransformer | 1M | 385K | 5M | 5M |
-| TimesNet | 1M | 1M | 151M | 2M |
-| PatchTST | 36M | 42M | 40M | 30M |
-| Crossformer | 43M | 43M | 18M | 216K |
+One notable advantage of DiPE-Linear is its substantially reduced parameter count compared to FC models.
+By operating within a low-dimensional manifold in the FC models’ model space, DiPE-Linear achieves performance comparable to or better than FC models, suggesting that our carefully designed model space provides a highly efficient and effective solution for these LTSF benchmark tasks.
+This highlights the robustness and effectiveness of our architecture in representing the essential solution space of FC models with minimal complexity.
+4.2.2 Versus nonlinear models.
+| DiPE-Linear | TimeMixer++ | TQNet | iTransformer | PatchTST | TimesNet | Crossformer | MICN | FEDformer | |||||||||||
+| (Ours) | (2025) | (2025) | (2024) | (2023) | (2023) | (2023) | (2023) | (2022) | |||||||||||
+| Metrics | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | MSE | MAE | |
+| ETTh1 | 96 | 0.369 | 0.393 | 0.361 | 0.403 | 0.371 | 0.393 | 0.386 | 0.405 | 0.370 | 0.399 | 0.384 | 0.402 | 0.386 | 0.429 | 0.396 | 0.427 | 0.376 | 0.415 |
+| 192 | 0.407 | 0.415 | 0.416 | 0.441 | 0.428 | 0.426 | 0.424 | 0.440 | 0.413 | 0.421 | 0.557 | 0.436 | 0.419 | 0.444 | 0.430 | 0.453 | 0.423 | 0.446 | |
+| 336 | 0.424 | 0.427 | 0.430 | 0.434 | 0.476 | 0.446 | 0.449 | 0.460 | 0.422 | 0.436 | 0.491 | 0.469 | 0.440 | 0.461 | 0.433 | 0.458 | 0.444 | 0.462 | |
+| 720 | 0.409 | 0.439 | 0.467 | 0.451 | 0.487 | 0.470 | 0.495 | 0.487 | 0.447 | 0.466 | 0.521 | 0.500 | 0.519 | 0.524 | 0.474 | 0.508 | 0.469 | 0.492 | |
+| ETTh2 | 96 | 0.275 | 0.336 | 0.276 | 0.328 | 0.295 | 0.343 | 0.297 | 0.348 | 0.274 | 0.336 | 0.340 | 0.374 | 0.276 | 0.338 | 0.289 | 0.357 | 0.332 | 0.374 |
+| 192 | 0.325 | 0.372 | 0.342 | 0.379 | 0.367 | 0.393 | 0.371 | 0.403 | 0.339 | 0.379 | 0.402 | 0.414 | 0.345 | 0.382 | 0.409 | 0.438 | 0.407 | 0.446 | |
+| 336 | 0.350 | 0.393 | 0.346 | 0.398 | 0.417 | 0.427 | 0.404 | 0.428 | 0.329 | 0.380 | 0.452 | 0.452 | 0.375 | 0.411 | 0.417 | 0.452 | 0.400 | 0.447 | |
+| 720 | 0.375 | 0.415 | 0.392 | 0.415 | 0.433 | 0.446 | 0.424 | 0.444 | 0.379 | 0.422 | 0.462 | 0.468 | 0.408 | 0.446 | 0.426 | 0.473 | 0.412 | 0.469 | |
+| ETTm1 | 96 | 0.309 | 0.350 | 0.310 | 0.334 | 0.311 | 0.353 | 0.300 | 0.353 | 0.290 | 0.342 | 0.338 | 0.375 | 0.316 | 0.373 | 0.314 | 0.360 | 0.326 | 0.390 |
+| 192 | 0.339 | 0.369 | 0.348 | 0.362 | 0.356 | 0.378 | 0.345 | 0.382 | 0.332 | 0.369 | 0.371 | 0.387 | 0.377 | 0.411 | 0.359 | 0.387 | 0.365 | 0.415 | |
+| 336 | 0.367 | 0.386 | 0.376 | 0.391 | 0.390 | 0.401 | 0.374 | 0.398 | 0.369 | 0.392 | 0.410 | 0.411 | 0.431 | 0.442 | 0.398 | 0.413 | 0.392 | 0.425 | |
+| 720 | 0.416 | 0.413 | 0.440 | 0.423 | 0.452 | 0.440 | 0.429 | 0.430 | 0.416 | 0.420 | 0.478 | 0.450 | 0.600 | 0.547 | 0.459 | 0.464 | 0.446 | 0.458 | |
+| ETTm2 | 96 | 0.162 | 0.252 | 0.170 | 0.245 | 0.173 | 0.256 | 0.175 | 0.266 | 0.165 | 0.255 | 0.187 | 0.267 | 0.421 | 0.461 | 0.178 | 0.273 | 0.180 | 0.271 |
+| 192 | 0.216 | 0.289 | 0.229 | 0.291 | 0.238 | 0.298 | 0.242 | 0.312 | 0.220 | 0.292 | 0.249 | 0.309 | 0.503 | 0.519 | 0.245 | 0.316 | 0.252 | 0.318 | |
+| 336 | 0.268 | 0.324 | 0.303 | 0.343 | 0.301 | 0.340 | 0.282 | 0.340 | 0.274 | 0.329 | 0.321 | 0.351 | 0.611 | 0.580 | 0.295 | 0.350 | 0.324 | 0.364 | |
+| 720 | 0.353 | 0.379 | 0.373 | 0.399 | 0.397 | 0.396 | 0.378 | 0.398 | 0.362 | 0.385 | 0.497 | 0.403 | 0.996 | 0.750 | 0.389 | 0.406 | 0.410 | 0.420 | |
+| Electricity | 96 | 0.132 | 0.228 | 0.135 | 0.222 | 0.134 | 0.229 | 0.132 | 0.229 | 0.129 | 0.222 | 0.168 | 0.272 | 0.187 | 0.283 | 0.159 | 0.267 | 0.186 | 0.302 |
+| 192 | 0.148 | 0.245 | 0.147 | 0.235 | 0.154 | 0.247 | 0.151 | 0.246 | 0.147 | 0.240 | 0.184 | 0.289 | 0.258 | 0.330 | 0.168 | 0.279 | 0.197 | 0.311 | |
+| 336 | 0.162 | 0.261 | 0.164 | 0.245 | 0.169 | 0.264 | 0.167 | 0.264 | 0.163 | 0.259 | 0.198 | 0.300 | 0.323 | 0.369 | 0.196 | 0.308 | 0.213 | 0.328 | |
+| 720 | 0.198 | 0.296 | 0.212 | 0.310 | 0.201 | 0.294 | 0.194 | 0.286 | 0.197 | 0.290 | 0.220 | 0.320 | 0.404 | 0.423 | 0.203 | 0.312 | 0.233 | 0.344 | |
+| Weather | 96 | 0.142 | 0.201 | 0.155 | 0.205 | 0.157 | 0.200 | 0.159 | 0.208 | 0.149 | 0.198 | 0.172 | 0.220 | 0.153 | 0.217 | 0.161 | 0.226 | 0.238 | 0.314 |
+| 192 | 0.187 | 0.253 | 0.201 | 0.245 | 0.206 | 0.245 | 0.200 | 0.248 | 0.194 | 0.241 | 0.219 | 0.261 | 0.197 | 0.269 | 0.220 | 0.283 | 0.275 | 0.329 | |
+| 336 | 0.234 | 0.293 | 0.237 | 0.265 | 0.262 | 0.287 | 0.253 | 0.289 | 0.245 | 0.282 | 0.280 | 0.306 | 0.252 | 0.311 | 0.275 | 0.328 | 0.339 | 0.377 | |
+| 720 | 0.306 | 0.348 | 0.312 | 0.334 | 0.344 | 0.342 | 0.321 | 0.338 | 0.314 | 0.334 | 0.365 | 0.359 | 0.318 | 0.363 | 0.311 | 0.356 | 0.389 | 0.409 | |
+| Best count | 15 | 11 | 2 | 9 | 0 | 1 | 1 | 1 | 8 | 5 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |
+Next, we compare DiPE-Linear with several nonlinear models.
+Table 2 presents the evaluation results for these nonlinear models specifically on the public LTSF datasets.
+We do not report model performance on the private LTSF and STSF datasets due to the following practical limitations: for private LTSF datasets, excessively long input lengths lead to high memory usage for nonlinear models, resulting in prohibitively high deployment costs in production environments.
+For STSF datasets, the limited amount of training data makes these models highly prone to overfitting, which complicates the training process.
+To ensure a rigorous comparison, baseline models were evaluated by selecting the best result from a comprehensive search over input lengths .
+To benchmark against the very latest advancements, we also include two recent models, TimeMixer++[18] and TQNet[9], citing the best performance reported in their original publications.
+The results in Table 2 highlight the remarkable effectiveness of DiPE-Linear.
+Notably, our model establishes a new state-of-the-art, achieving the best overall accuracy against all competing methods.
+It significantly outperforms established baselines like iTransformer and recent models such as TimeMixer++[18] and TQNet[9].
+This SOTA performance is achieved with unparalleled efficiency.
+For instance, DiPE-Linear matches or exceeds the accuracy of PatchTST [3] while utilizing only of its parameters.
+This massive reduction in model complexity, combined with its superior forecasting accuracy, underscores the competitive advantage of DiPE-Linear and establishes it as a new, highly efficient benchmark for LTSF.
+Furthermore, our state-of-the-art results, achieved with a purely linear model, contribute strong evidence to the ongoing debate [23] regarding the necessity of complex, non-linear architectures for LTSF.
+The success of DiPE-Linear suggests that, for these benchmarks, a well-structured inductive bias (i.e., the factorization of frequential and temporal components) is more critical for generalization than the capacity for high-dimensional non-linear mapping (as found in Transformers).
+4.2.3 Parameter complexity and efficiency.
+Recent studies have focused on parameter efficient models.
+We compare DiPE-Linear with SparseTSF [10] and FITS [21] (FITS-T and FITS-L, representing different low-pass filter settings).
+As shown in Fig.
+4, our model achieves the lowest MSE across these parameter-efficient baselines.
+Notably, DiPE-Linear not only uses fewer parameters than FITS-T but also matches or outperforms FITS-L, demonstrating a superior balance of performance and parameter efficiency.
 Fig.
@@
 The results highlight that DiPE-Linear achieves state-of-the-art performance while maintaining a drastically smaller parameter count, underscoring its efficiency and practicality.
-IV-B4 Effect of Training Data Size
+4.2.4 Effect of Training Data Size.
 To further investigate whether the model can perform effectively with reduced training data, we conducted experiments on the ETTh1 and ETTm1 datasets using different proportions of the training set.
-For the ETTh1 dataset, we compared our model with other FCs.
+For the ETTh1 dataset, we compared our model with other FC models.
 Additionally, since the ETTm1 dataset contains more samples, we included nonlinear models for a broader comparison.
 As shown in Fig.
-6, DiPE-Linear outperforms both FCs and nonlinear models across all proportions of the training set.
+5, DiPE-Linear outperforms both FC models and nonlinear models across all proportions of the training set.
 Notably, on the ETTh1 dataset, our model surpasses the performance of other FC models trained on the full dataset while using only 40% of the training data.
-IV-C Ablation Studies
-IV-C1 Basic Components
+4.3 Ablation Studies
+4.3.1 Basic Components.
 To comprehensively validate the effectiveness of our model, we conducted ablation tests on two submodules: SFA and STA.
 The submodule IFM, which maps the input sequence space to the output sequence space, serves as the core component of our model and therefore cannot be ablated.
-Table V presents the results of the ablation experiments on the ETTh2 and Weather datasets with various prediction lengths.
+Table 3 presents the results of the ablation experiments on the ETTh2 and Weather datasets with various prediction lengths.
 In both the ETTh2 and Weather datasets, the incorporation of the SFA and STA modules generally enhances the model’s prediction accuracy.
@@
 | SFA | STA | IFM | 96 | 192 | 336 | 720 | 96 | 192 | 336 | 720 |
-| ✗ | ✓ | ✓ | 0.284 | 0.331 | 0.348 | 0.382 | 0.146 | 0.191 | 0.243 | 0.315 |
-| ✓ | ✗ | ✓ | 0.290 | 0.331 | 0.348 | 0.379 | 0.166 | 0.220 | 0.272 | 0.331 |
-| ✗ | ✗ | ✓ | 0.295 | 0.340 | 0.352 | 0.387 | 0.175 | 0.228 | 0.274 | 0.332 |
+| ✓ | 0.295 | 0.340 | 0.352 | 0.387 | 0.175 | 0.228 | 0.274 | 0.332 | ||
+| ✓ | ✓ | 0.284 | 0.331 | 0.348 | 0.382 | 0.146 | 0.191 | 0.243 | 0.315 | |
+| ✓ | ✓ | 0.290 | 0.331 | 0.348 | 0.379 | 0.166 | 0.220 | 0.272 | 0.331 | |
 | ✓ | ✓ | ✓ | 0.275 | 0.325 | 0.350 | 0.375 | 0.142 | 0.187 | 0.234 | 0.306 |
-IV-C2 Low-rank weight sharing
-In multivariate time series, our model must effectively capture the relationships between variables.
-We adopt a low-rank weight sharing approach, specifying the learning of sets of weights.
-Figure 7 presents the results of testing different values of on the Weather dataset with a prediction length of .
-When , our model is equivalent to weight sharing, meaning all variables use the same set of weights.
-When , our model learns a separate set of weights for each variable, treating the variables as independent.
-The experiments indicate that selecting an appropriate value for enables the model to achieve optimal performance.
-We conducted an in-depth analysis to evaluate the necessity of low-rank weight sharing for each component of the model.
-Extensive experiments were performed on the ETTh2 and Weather datasets, covering multiple prediction horizons.
-For each component, we compared three configurations: shared weights, low-rank shared weights, and full-rank independent weights.
-As presented in Table VI, the low-rank weight sharing consistently delivered superior performance across all settings, highlighting its critical role in optimizing model efficacy.
-| Components | ETTh2 | Weather | ||||||||
-| SFA | STA | IFM | 96 | 192 | 336 | 720 | 96 | 192 | 336 | 720 |
-| ✗ | ✓ | ✓ | 0.280 | 0.329 | 0.353 | 0.375 | 0.143 | 0.187 | 0.238 | 0.310 |
-| ✓ | ✗ | ✓ | 0.279 | 0.326 | 0.354 | 0.374 | 0.142 | 0.187 | 0.238 | 0.309 |
-| ✓ | ✓ | ✗ | 0.283 | 0.329 | 0.353 | 0.378 | 0.142 | 0.188 | 0.238 | 0.310 |
-| F | F | F | 0.270 | 0.333 | 0.356 | 0.380 | 0.169 | 0.214 | 0.276 | 0.340 |
-| ✓ | ✓ | ✓ | 0.275 | 0.325 | 0.350 | 0.375 | 0.142 | 0.187 | 0.234 | 0.306 |
-IV-C3 SFALoss
-The parameter in SFALoss governs the trade-off between the frequency domain loss and the time domain loss.
-To assess the influence of on model performance, we conducted experiments on the ETTh1 and ETTh2 datasets with a prediction length of .
-The results, illustrated in Figure 8, demonstrate the significant impact of on optimizing the model’s predictive accuracy across both datasets.
-IV-D Interpretability: Case Studies
-Benefiting from the disentangled representation of our model in both the frequency and time domains, our approach demonstrates superior interpretability.
-This interpretability facilitates user understanding of the model’s behavior, thereby enhancing trust in its predictive outcomes.
-IV-D1 ETTh1
-We select the ETTh1 dataset as a simple example.
-We choose an input length and a prediction horizon , ensuring the model has the minimum number of parameters.
-Additionally, within the ETTh1 dataset, we set the hyperparameter , which eliminates the router .
-Thus for each module, we learn only a single set of weights.
-The weight visualization is shown in Fig.
-9.
-The SFA module explicitly preserves low-frequency information while emphasizing harmonic amplitudes in the high-frequency region.
-The STA module identifies the importance of each time point in the input sequence for future predictions.
-The IFM module learns distinct prediction patterns for harmonic frequencies (shown in red) and non-harmonic frequencies (shown in blue).
-IV-D2 Weather
-In the Weather dataset, different variables represent various meteorological data, such as temperature, pressure, humidity, wind direction, precipitation, etc..
-We set as the hyperparameter, allowing the model to learn that the router can represent the correlations among different meteorological data variation patterns.
-The visualization is shown in Fig.
-10.
-We can further process the weights learned by the model to understand the knowledge it has captured.
-Taking the Router as an example, for each variable, the router weight can be viewed as the distribution of its predictive patterns across different weight groups.
-Consequently, we can compute the Jensen-Shannon Distance (JSD) between variables:
-where are two variables, is a mixture distribution, and is the Kullback–Leibler (KL) divergence defined by 17:
-Thus, we obtained the JSD Matrix based on the learnt router, as shown in Fig. 11.
-IV-D3 Electricity
-We further selected the Electricity dataset to comprehensively demonstrate the representations learned by DiPE-Linear.
-We chose an input length and a prediction horizon in the LSTF setting.
-In the Electricity dataset, the hyperparameter , meaning that four sets of weights are learned for each module.
-The parameter visualization is shown in Figure 10.
-Similar to the ETTh1 dataset, the SFA module learns to retain low-frequency information and high-frequency harmonics.
-The STA module clearly distinguishes the periodic variations between weekdays and weekends within a month.
-In the IFM module, four different frequency transformation patterns are also learned.
-Finally, the weight visualization of the Router illustrates the differences and similarities in predictive patterns across different variables in the dataset.
-IV-E Complexity analysis
-Our model retains the linear characteristics of FCs while achieving significant reductions in both parameter and computational complexity.
-In this section, we analyze the complexity of each module in our model in detail.
-For simplicity, we denote ambiguously both the input length and forecast horizon as .
-IV-E1 Parameter Complexity
-IV-E2 Computational Complexity
-The computational cost of our model primarily arises from two components.
-First, the model’s parameter computations involve only element-wise multiplications, yielding a computational complexity of .
-Second, the rFFT and irFFT operations each have a computational complexity of .
-Therefore, the overall computational complexity of the model is .
-A comparison of the parameter and computational complexities between our model and FCs (both with and without weight sharing) is provided in Table VII.
+4.3.2 Hyperparameter Analysis and Guidelines.
+We analyze the sensitivity of the two key hyperparameters, and , to provide practical tuning guidelines and address concerns about model complexity.
+For the Low-rank weight sharing factor , this hyperparameter can be intuitively understood as the number of distinct patterns among the variables in the multivariate time series.
+Fig.
+6a presents the results of varying on the Weather dataset ().
+The key observation is that model performance remains stable and near-optimal across a wide range of (from 2 to 16).
+Performance only degrades at the extremes: (full weight sharing) and (full channel independence).
+This demonstrates that is not a sensitive hyperparameter within a reasonable range, significantly easing the tuning process.
+In practice, we suggest starting and increasing it exponentially(as ) until validation performance saturates.
+For the SFALoss balance parameter , which governs the trade-off between the time-domain loss () and the frequency-domain loss (), we performed a grid search on the ETTh datasets (results in Fig.
+6).
+Our experiments show that the model’s accuracy often has a monotonic or U-shaped relationship with .
+Notably, the optimal performance is typically found when is in the ranges of [0, 0.3] or [0.7, 1.0].
+This suggests the model prefers to optimize primarily for one domain (either time or frequency) rather than an equal mixture.
+This finding simplifies tuning, as the search space can be effectively concentrated on values near 0 and 1, rather than requiring a fine-grained search across the entire [0, 1] interval.
+4.4 Complexity Analysis
+Our model retains the linear parameterization of FC models while significantly reducing both parameter and computational complexity.
+The parameter complexity in the SFA, STA, and IFM modules is .
+Low-rank weight sharing introduces additional routing parameters with a complexity of , where is the number of channels; this is a higher-order infinitesimal relative to .
+Thus, the overall parameter complexity remains .
+For computational complexity, the model’s main costs arise from element-wise multiplications () and the rFFT/irFFT operations (), resulting in a total computational complexity of .
+Table 4 compares these complexities with those of standard FC models, demonstrating the efficiency of our approach.
+We note that while our theoretical computational complexity is superior, practical latency depends on hardware optimizations; FC models benefit from highly optimized matrix multiplication (e.g.
+cuBLAS), which may not be true for FFT operations.
+Therefore, the primary advantages of DiPE-Linear are its linear parameter complexity () and drastically reduced memory footprint.
+This makes DiPE-Linear ideal for resource-constrained environments and for mitigating overfitting on long sequences.
 | DiPE-Linear (Ours) | FC (w/ sharing) | FC (w/o sharing) | |
-| Param. | |||
-| Compu. |
-V Conclusion
-This paper presents DiPE-Linear, a novel approach designed for LTSF in industrial applications.
-DiPE-Linear effectively addresses the challenges of Industry 5.0, including high-dimensional and high-resolution data as well as high-stakes scenarios, by leveraging disentangled, linear-complexity interpretable modules in both the time and frequency domains.
-The sophisticated design reduces parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear.
-Extensive experimental results demonstrate that DiPE-Linear consistently outperforms existing models, significantly enhancing scalability, computational efficiency, and predictive accuracy, particularly in complex industrial environments.
-In future work, we will enhance DiPE-Linear by incorporating interpretable nonlinear modules, while further exploring lightweight and efficient practical LSTF models.
+| Parameter | |||
+| Time |
+5 Conclusion
+This paper introduces DiPE-Linear, a novel model specifically tailored for LTSF.
+DiPE-Linear addresses the intrinsic quadratic parameter redundancy and entanglement of existing monolithic FC models by employing disentangled modules—factorizing the modeling task into specialized components for frequential filtering, temporal weighting, and frequential mapping.
+This innovative architecture, which serves as a strong inductive bias, reduces parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear, significantly improving efficiency and scalability.
+Comprehensive experimental evaluations demonstrate that DiPE-Linear consistently achieves or surpasses state-of-the-art models, suggesting that for LTSF benchmarks, its well-structured linear design is more effective for generalization than the complex non-linear mapping of larger models.
+DiPE-Linear thus sets a new baseline for parameter-efficient LTSF models.
+Future research will explore the integration of nonlinear architectures and further optimization of efficient LTSF frameworks.